### v1

# Rectangular recurrence relations in $\mathfrak{gl}_{n}$ and $\mathfrak{o}_{2n+1}$ invariant integrable models

Abstract.
A new method is introduced to derive general recurrence relations for off-shell Bethe vectors in quantum integrable models with either type $\mathfrak{gl}_{n}$ or type $\mathfrak{o}_{2n+1}$ symmetries. These recurrence relations describe how to add a single parameter, $z$, to specific subsets of Bethe parameters, expressing the resulting Bethe vector as a linear combination of monodromy matrix entries that act on Bethe vectors which do not depend on $z$. We refer to these recurrence relations as rectangular because the monodromy matrix entries involved are drawn from the upper-right rectangular part of the matrix. This construction is achieved within the framework of the zero mode method.
1 Introduction
We will consider generic $\mathfrak{g}$-invariant quantum integrable models in the framework of the algebraic Bethe ansatz [2]. In such models the monodromy matrix depends on a spectral parameter and satisfies the $RTT$ relations [3] with a $\mathfrak{g}$-invariant $R$-matrix. In the Hilbert space of physical states, one can always construct from the local operators of the model a basis of states which form a representation of the finite-dimensional algebras $\mathfrak{g}$ and that are eigenvectors of a set of commuting Hamiltonians. In the algebraic Bethe ansatz these states are constructed from the monodromy matrix entries which depend on spectral parameters satisfying the Bethe equations. These states are called on-shell Bethe vectors. When the spectral parameters are generic (not forced to obey the Bethe equations), the Bethe vectors are called off-shell and their combinatorial properties are defined solely by the $RTT$ relation with a given $R$-matrix. As a consequence, one can replace the monodromy matrix of a generic model by the fundamental $T$-operator of the Yangian $Y(\mathfrak{g})$ in its matrix realization [1, 23]. The commutation relation for this $T$-operator $T(u)$, which depends on a formal spectral parameter $u$, coincides with the commutation relations of monodromy matrix in a generic $\mathfrak{g}$-invariant integrable model and the generators of the finite dimensional symmetry may be identified with the zero modes of the matrix entries $T_{i,j}(u)$ of the Yangian fundamental $T$-operator. Below we will explore the Yangian fundamental $T$-operator $T(u)$ calling it the monodromy matrix of a generic $\mathfrak{g}$-invariant integrable model.
One of the fundamental problems in quantum integrable models is the presentation of the Bethe vectors scalar products in a determinant form. For periodic boundary conditions, this problem was investigated and fully solved in [25] for the integrable system associated with the simplest Lie algebra $\mathfrak{gl}_{2}$, where the structure of the Bethe vectors is quite simple.
In the models associated to higher rank symmetries the structure of the Bethe vectors is rather complicated. In principle, the nested Bethe ansatz ensures that the Bethe vectors can be expressed as a combinatorial expression of the monodromy matrix entries acting on a vacuum vector. However, no reasonable formulas for the scalar products can be obtained from this combinatorial presentation.
On the other hand, if one can find formulas for the action of the monodromy matrix entries on the off-shell Bethe vectors, as well as recurrence relations for them, then one can find expressions for the scalar products of off-shell Bethe vectors, and ultimately obtain recurrence relations for the building blocks of these scalar products. Expressions for the scalar product and the norm of Bethe vectors were obtained in [12, 10] for $\mathfrak{gl}_{2}$ invariant models and in [24] for $\mathfrak{gl}_{3}$ invariant models. General expressions for the scalar product, the norm and the recurrence relations were achieved in [4, 5] for models associated to $\mathfrak{gl}(m|n)$. Here by recurrence relations for the off-shell Bethe vectors we mean the possibility to express a Bethe vector with an extended set of Bethe parameters (for example, with an extension by a parameter $z$) as a linear combination of the action of monodromy matrix entries $T_{i,j}(z)$ with $i<j$ acting on off-shell Bethe vectors which do not depend on $z$.
The action formulas of the monodromy entries on the off-shell Bethe vectors in $\mathfrak{gl}(m|n)$- and $\mathfrak{o}_{2n+1}$-invariant integrable models were given in [7] and [18] respectively. The existence of recurrence relations is less investigated. Some examples of such relations can be found in the aforementioned papers. One of the goal of this paper is to fill this gap and produce all possible recurrence relations for the Bethe vectors in the models with $\mathfrak{gl}_{n}$ and $\mathfrak{o}_{2n+1}$ symmetries.
In the present paper, we show that in order to solve this problem, one has to combine a single (simple) action of the monodromy matrix entry with the action of the zero modes of the monodromy matrix, the latter being identified with the simple root generators of the Lie algebra $\mathfrak{g}$. We first perform the calculation of the recurrence relations using this method for the Bethe vectors in $\mathfrak{gl}_{n}$-invariant models and then extend this approach to the Bethe vectors in $\mathfrak{o}_{2n+1}$-invariant models. The obtained recurrence relations will be then tested in several limiting cases to observe the reductions over rank of $\mathfrak{g}$ and embeddings of $Y(\mathfrak{gl}_{\ell})\otimes Y(\mathfrak{gl}_{n-\ell})$ into $Y(\mathfrak{gl}_{n})$ and of $Y(\mathfrak{o}_{2\ell+1})\otimes Y(\mathfrak{gl}_{n-\ell})$ into $Y(\mathfrak{o}_{2n+1})$.
The plan of the paper is as follows. In section 2, we recall some basic algebraic notions on Yangians, which are at the core of the models under consideration, as well as some definitions on Bethe vectors. Section 3 deals with the rectangular recurrence relations for $\mathfrak{gl}_{n}$ models. Here we verify also the consistency of the recurrence relations with the embedding $Y(\mathfrak{gl}_{\ell})\otimes Y(\mathfrak{gl}_{n-\ell})$ into $Y(\mathfrak{gl}_{n})$ and compare our approach with the recent papers [13, 14]. The case of $\mathfrak{o}_{2n+1}$ is studied in section 4. The main result is contained in section 4.1, where rectangular recurrence relations for $\mathfrak{o}_{2n+1}$ models are presented. Some particular cases and examples are displayed in section 4.2. We also show that the recurrence relations are consistent with the embedding of $Y(\mathfrak{gl}_{n})$ or $Y(\mathfrak{o}_{2a+1})\otimes Y(\mathfrak{gl}_{n-a})$ in $Y(\mathfrak{o}_{2n+1})$: the Bethe vectors exhibit a nice factorisation property over these subalgebras (sections 4.3 and 4.4). We conclude in section 5, and two appendices are devoted to the technical proofs.
2 Preliminaries
2.1 Algebraic context
Let $\mathfrak{g}$ be either the classical Lie algebra $\mathfrak{gl}_{n}$ or the orthogonal algebra $\mathfrak{o}_{2n+1}$, where $n=2,3,\ldots$. We will use the set of positive integers $I_{\mathfrak{gl}_{n}}=\{1,\ldots,n\}$ to index elements of the operators in End$({{\mathbb{C}}}^{n})$ and the set of integers $I_{\mathfrak{o}_{2n+1}}=\{-n,-n+1,\ldots,-1,0,1,2,\ldots,n\}$ to index elements of the operators in End$({{\mathbb{C}}}^{2n+1})$. We will use the notation $I_{\mathfrak{g}}$ to describe the two sets of indices simultaneously. Let $N=n$ and $N=2n+1$ for the algebras $\mathfrak{gl}_{n}$ and $\mathfrak{o}_{2n+1}$ respectively.
$RTT$ presentation of the Yangians $Y(\mathfrak{gl}_{n})$ and $Y(\mathfrak{o}_{2n+1})$.
Let $R_{\mathfrak{g}}(u,v)$ be the $\mathfrak{g}$-invariant $R$-matrix [26, 27]
$R_{\mathfrak{g}}(u,v)=\mathbf{I}\otimes\mathbf{I}+\frac{c\,\mathbf{P}}{u-v}-\frac{c\,\mathbf{Q}}{u-v+c\kappa_{n}}\,,$ (2.1)
where $\mathbf{I}=\sum_{i\in I_{\mathfrak{g}}}{\sf e}_{i,i}$ is the identity operator acting in the space ${\mathbb{C}}^{N}$ and ${\sf e}_{i,j}\in{\rm End}({\mathbb{C}}^{N})$ are $N\times N$ matrices with the only nonzero entry equals to 1 at the intersection of the $i$-th row and $j$-th column. The operators $\mathbf{P}$ and $\mathbf{Q}$ act in ${\mathbb{C}}^{N}\otimes{\mathbb{C}}^{N}$. They read
$\mathbf{P}=\sum_{i,j\in I_{\mathfrak{g}}}{\sf e}_{i,j}\otimes{\sf e}_{j,i},\quad\mathbf{Q}=\sum_{i,j\in I_{\mathfrak{g}}}{\sf e}_{-i,-j}\otimes{\sf e}_{i,j}\,,$
and [^1] The value of $\kappa_{n}=\infty$ for the algebra $\mathfrak{gl}_{n}$ simple means that in this case the $\mathfrak{g}$-invariant $R$-matrix does not contain the term proportional to operator $\mathbf{Q}$ and so coincides with the Yang $R$-matrix [26]. The parameter $\kappa_{n}=n-1/2$ is relevant only for the algebra $\mathfrak{g}=\mathfrak{o}_{2n+1}$.
$\kappa_{n}=\begin{cases}\infty\quad&{\rm for}\quad\mathfrak{g}=\mathfrak{gl}_{n}\,,\\ n-1/2\quad&{\rm for}\quad\mathfrak{g}=\mathfrak{o}_{2n+1}\,.\end{cases}$ (2.2)
For any matrix $M\in{\rm End}({\mathbb{C}}^{2n+1})$ we denote by $M^{\rm t}$ the transposition with respect to the secondary diagonal
$\big{(}M^{\rm t}\big{)}_{i,j}=M_{-j,-i}\,.$ (2.3)
In particular, $\mathbf{Q}=\mathbf{P}^{{\rm t}_{1}}=\mathbf{P}^{{\rm t}_{2}}$.
The generic $\mathfrak{g}$-invariant integrable model is described by the monodromy matrix $T(u)$ which depends on the spectral parameter $u$ and satisfies the commutation relation
$R_{\mathfrak{g}}(u,v)\left(T(u)\otimes\mathbf{I}\right)\left(\mathbf{I}\otimes T(v)\right)=\left(\mathbf{I}\otimes T(v)\right)\left(T(u)\otimes\mathbf{I}\right)R_{\mathfrak{g}}(u,v),$ (2.4)
with the $\mathfrak{g}$-invariant $R$-matrix (2.1).
Equation (2.4) yields the commutation relations of the monodromy matrix entries
$\begin{split}&\left[T_{i,j}(u),T_{k,l}(v)\right]=\frac{c}{u-v}\left(T_{k,j}(v)T_{i,l}(u)-T_{k,j}(u)T_{i,l}(v)\right)\\ &\quad+\frac{c}{u-v+c\kappa_{n}}\sum_{p=-n}^{n}\left(\delta_{k,-i}\,\,T_{p,j}(u)T_{-p,l}(v)-\delta_{l,-j}\,T_{k,-p}(v)T_{i,p}(u)\right).\end{split}$ (2.5)
Note that the second line in (2.5) occurs only in the case of $\mathfrak{o}_{2n+1}$ models, in accordance with the footnote 1. We will assume the following dependence of the monodromy matrix on the spectral parameter
$T_{i,j}(u)=\chi_{i}\ \delta_{ij}+\sum_{\ell\geq 0}T_{i,j}[\ell](u/c)^{-\ell-1}\,,$ (2.6)
where the parameters $\chi_{i}$, $i\in I_{\mathfrak{g}}$ are twisting parameters. For $\mathfrak{g}=\mathfrak{gl}_{N}$ these parameters are all independent, while for $\mathfrak{g}=\mathfrak{o}_{2n+1}$ they satisfy the relation $\chi_{i}\chi_{-i}=1$ and $\chi_{0}=1$. For both algebras, the possibility to start the series expansion (2.6) with $\chi_{i}\,\delta_{ij}$ instead of $\delta_{ij}$ corresponds to the usual twist $T(z)\to D\,T(z)$ where $D$ is the diagonal matrix with $D_{ii}=\chi_{i}$. These twisting maps are isomorphisms of the Yangians, since the $R$-matrix commutes with $D\otimes D$ (with the restriction $D^{\rm t}=D$ in the $\mathfrak{o}_{2n+1}$ case).
Central element for $\mathfrak{g}=\mathfrak{o}_{2n+1}$. Remark that for $\mathfrak{g}=\mathfrak{o}_{2n+1}$, the pole at $u=v-c\kappa_{n}$ in the $R$-matrix (2.1) and in the commutation relation (2.4) implies that the monodromy matrix should satisfy following relations [11]
$T(z)^{\rm t}\cdot T(z+c\kappa_{n})=T(z+c\kappa_{n})\cdot T(z)^{\rm t}=\mathcal{C}(z)\ \mathbf{I}\,,$ (2.7)
where $\mathcal{C}(z)$ is a central operator in the algebra defined by the relations (2.4). Let $\pi(z)$ be a formal series $\pi(z)=1+\sum_{m\geq 0}\pi_{m}\,(c/z)^{-m-1}$ which solves the equation
$\mathcal{C}(z)=\pi(z)\,\pi(z+c\kappa_{n})\,.$ (2.8)
This equation can be solved inductively expressing the coefficient of the formal series $\pi(z)$ through the coefficients of the central element $\mathcal{C}(z)$. Then one can rescale the monodromy matrix
$T(z)\to T^{\prime}(z)=\pi(z)^{-1}\,T(z)$
in such a way that the rescaled monodromy matrix satisfies the equation
$T^{\prime}(z)^{\rm t}\cdot T^{\prime}(z+c\kappa_{n})=T^{\prime}(z+c\kappa_{n})\cdot T^{\prime}(z)^{\rm t}=\mathbf{I}\,.$ (2.9)
Zero modes.
The zero modes $T_{i,j}[0]\equiv{\sf T}_{i,j}$ will play an important role in our approach. For the spectral parameter dependence as in (2.6) they are defined as the coefficients at $c/u$. More generally, in each integrable models, the monodromy matrix zero mode operators may be always defined as the operators built from the monodromy matrix entries which satisfy the commutation relations of the finite-dimensional algebra $\mathfrak{g}$.
Using the definition of the zero modes
${\sf T}_{i,j}=\mathop{\rm lim}_{u\to\infty}\frac{u}{c}\ (T_{i,j}(u)-\chi_{i}\delta_{i,j})$ (2.10)
one gets from (2.5) the commutation relations
$[{\sf T}_{i,j},T_{k,l}(u)]=\chi_{i}\,\delta_{i,l}\,{\sf T}_{k,j}(u)-\chi_{j}\,\delta_{k,j}\,T_{i,l}(u)$ (2.11)
for $\mathfrak{g}=\mathfrak{gl}_{n}$ and
$[{\sf T}_{i,j},T_{k,l}(u)]=\chi_{i}\,(\delta_{i,l}\,T_{k,j}(u)-\delta_{l,-j}\,T_{k,-i}(u))-\chi_{j}\,(\delta_{k,j}\,T_{i,l}(u)-\delta_{k,-i}\,T_{-j,l}(u))$ (2.12)
for $\mathfrak{g}=\mathfrak{o}_{2n+1}$.
Embeddings.
We describe different embeddings in the Yangians $Y(\mathfrak{gl}_{n})$ and $Y(\mathfrak{o}_{2n+1})$ that will be reflected in a factorisation property of the Bethe vectors, see sections 4.3 and 4.4 below.
Embedding $Y(\mathfrak{gl}_{a})\otimes Y(\mathfrak{gl}_{n-a})\hookrightarrow Y(\mathfrak{gl}_{n})$. From the commutation relations (2.5), it is clear that in $Y(\mathfrak{gl}_{n})$ the elements $T_{i,j}(z)$, $1\leq i,j\leq a$ generate a Yangian subalgebra $Y(\mathfrak{gl}_{a})$ in $Y(\mathfrak{gl}_{n})$, while the elements $T_{a+i,a+j}(z)$, $1\leq i,j\leq n-a$ generate the Yangian subalgebra $Y(\mathfrak{gl}_{n-a})$. However, these two subalgebras do not commute. To get an embedding of $Y(\mathfrak{gl}_{a})\otimes Y(\mathfrak{gl}_{n-a})$ in $Y(\mathfrak{gl}_{n})$, one needs to consider quantum minors, see relation (1.84), corollary 1.7.2 and corollary 1.11.4 in [23].
Embeddings of $Y(\mathfrak{gl}_{n})\hookrightarrow Y(\mathfrak{o}_{2n+1})$. There are several ways to embed $Y(\mathfrak{gl}_{n})$ in $Y(\mathfrak{o}_{2n+1})$. We describe here four of them that have some relevance for the study of Bethe vectors. Let us denote the $Y(\mathfrak{gl}_{n})$ monodromy matrix embedded in $Y(\mathfrak{o}_{2n+1})$ as ${\mathbb{T}}(z)$.
  • From the commutation relations (2.5), it is clear that in $Y(\mathfrak{o}_{2n+1})$ the elements
    $T_{i,j}(z)\equiv{\mathbb{T}}_{i,j}(z)\,,\quad 1\leq i,j\leq n$ (2.13)
    generate the Yangian $Y(\mathfrak{gl}_{n})$, since they obey the commutation relations (2.5) without the second line (all the terms which include the Kronecker symbol $\delta$ vanish).
  • Another embedding can be done, considering the elements $T_{i,j}(z)$, $-n\leq i,j\leq-1$ and defining ${\mathbb{T}}_{i,j}(z)=T_{-i,-j}(z)$, $1\leq i,j\leq n$. It is easy to see that ${\mathbb{T}}_{i,j}(z)$ also obey the commutation relations (2.5) without the second line. The twisting parameters for the monodromy ${\mathbb{T}}(z)$ in this case are $\chi_{i}^{-1}$.
  • A third embedding can be constructed through the relation (2.9). Indeed, from the relation ${T}(z)=\big{(}T(z+c\kappa_{n})^{-1}\big{)}^{\rm t}$, one deduces that the generators ${\mathbb{T}}_{i,j}(z)=\big{(}T(z+c\kappa_{n})^{-1}\big{)}_{-j,-i}$, for $1\leq i,j\leq n$, also obey the $Y(\mathfrak{gl}_{n})$ commutation relations.
Finally, a fourth embedding can be constructed, that will be used below.
Lemma 2.1.
The monodromy matrix ${\mathbb{T}}(z)$ with matrix entries given by
${\mathbb{T}}_{{i^{\prime}},{j^{\prime}}}(z)=-T_{i,j}(z)\,(-1)^{\delta_{i,j}+{{\sf T}_{i,i}/\chi_{i}+{\sf T}_{j,j}/\chi_{j}}},\ \text{with }\begin{cases}{i^{\prime}}=i+n+1,\\ {j^{\prime}}=j+n+1,\end{cases}-n\leq i,j\leq-1\,,$ (2.14)
satisfies the $\mathfrak{gl}_{n}$-type $RTT$ commutation relation with $R$-matrix
$R_{\mathfrak{gl}_{n}}(u,v)=\mathbf{I}\otimes\mathbf{I}+\frac{c\,\mathbf{P}}{u-v}\,.$ (2.15)
Proof.
The monodromy entries $T_{i,j}(u)$ for $-n\leq i,j\leq-1$ of the original $\mathfrak{o}_{2n+1}$ monodromy $T(u)$ satisfy the $\mathfrak{gl}_{n}$-type commutation relations which means that
$[T_{i,j}(u),T_{k,\ell}(v)]=g(u,v)\Big{(}T_{k,j}(v)\ T_{i,\ell}(u)-T_{k,j}(u)\ T_{i,\ell}(v)\Big{)}$
and
$[{{\sf T}_{i,i}},T_{k,\ell}(v)]=\chi_{i}\,(\delta_{i,l}-\delta_{i,k})\ T_{k,\ell}(v)\,.$
The latter equation can be written in the form
$(-1)^{{\sf T}_{i,i}/\chi_{i}}\ T_{k,\ell}(v)\ (-1)^{-{\sf T}_{i,i}/\chi_{i}}=(-1)^{\delta_{i,\ell}}\ (-1)^{\delta_{i,k}}\ T_{k,\ell}(v)\,.$
Using this equality one can verify that the monodromy entries ${\mathbb{T}}_{{i^{\prime}},{j^{\prime}}}(u)$ satisfy the $\mathfrak{gl}_{n}$-type commutation relation
$[{\mathbb{T}}_{{i^{\prime}},{j^{\prime}}}(u),{\mathbb{T}}_{{k^{\prime}},{\ell^{\prime}}}(v)]=g(u,v)\Big{(}{\mathbb{T}}_{{k^{\prime}},{j^{\prime}}}(v)\ {\mathbb{T}}_{{i^{\prime}},{\ell^{\prime}}}(u)-{\mathbb{T}}_{{k^{\prime}},{j^{\prime}}}(u)\ {\mathbb{T}}_{{i^{\prime}},{\ell^{\prime}}}(v)\Big{)}\,.$
$\square$
Embedding $Y(\mathfrak{o}_{2a+1})\otimes Y(\mathfrak{gl}_{n-a})\hookrightarrow Y(\mathfrak{o}_{2n+1})$. This type of embedding is more intricate. It uses the concept of quasi-determinants [22] $\widehat{T}_{i,j}(z)$ of the monodromy matrix, and we refer to e.g. [23] for more details about quasi-determinants. Indeed, one can show that $\widehat{T}_{i,j}(z)$, $1\leq i,j\leq a$ generate a Yangian subalgebra $Y(\mathfrak{o}_{2a+1})$, while $T_{k+a,\ell+a}(z)$, $1\leq k,\ell\leq n-a$ generate a Yangian subalgebra $Y(\mathfrak{gl}_{n-a})$. Moreover, we have $[\widehat{T}_{i,j}(z)\,,\,T_{k+a,\ell+a}(z)]=0$ for $1\leq i,j\leq a$ and $1\leq k,\ell\leq n-a$. A detailed presentation of this approach is beyond the scope of the present article, we refer to theorem 3.7 and corollary 3.10 in [11] for a detailed presentation of the construction. It is remarkable that this type of embedding can be seen in a simple way on the structure of Bethe vectors, see section 4.4 below.
2.2 Bethe vectors
In the framework of algebraic Bethe ansatz, the states in the Hilbert space of the physical model are defined by the vectors $|0\rangle$ and $\langle 0|$ such that
$T_{i,j}(z)|0\rangle=0\,,\quad i>j\,,\quad T_{i,i}(z)|0\rangle=\lambda_{i}(z)|0\rangle\,,\quad i\in I_{\mathfrak{g}}$ (2.16)
and
$\langle 0|T_{i,j}(z)=0\,,\quad i<j\,,\quad\langle 0|T_{i,i}(z)=\lambda_{i}(z)\langle 0|\,,\quad i\in I_{\mathfrak{g}}\,.$ (2.17)
In (2.16), the monodromy matrix elements are acting to the right, while in (2.17) they are acting to the left. If they exist, such vectors are called vacuum vectors.
The functions $\lambda_{i}(z)$ are characterizing the physical model under consideration. Since we are considering generic model we will consider these functions as free functional parameters. For $\mathfrak{gl}_{n}$-invariant integrable models, the functions $\lambda_{i}(z)$ are all independent, while for $\mathfrak{o}_{2n+1}$-invariant models and due to (2.9) they satisfy the relations [17]
$\lambda_{-j}(z)=\frac{1}{\lambda_{j}(z_{j})}\ \prod_{s=j+1}^{n}\frac{\lambda_{s}(z_{s-1})}{\lambda_{s}(z_{s})},\qquad j=0,1,\ldots,n\,,$ (2.18)
where we introduced the shifted spectral parameter
$z_{s}=z-c\Big{(}s-\frac{1}{2}\Big{)},\qquad s=0,1,\ldots,n\,.$ (2.19)
Due to our choice (2.6) of dependence of monodromy matrix entries on the formal spectral parameter the free functional parameters $\lambda_{i}(z)$ are formal series
$\lambda_{i}(z)=\chi_{i}+\sum_{\ell\geq 0}\lambda_{i}[\ell]\ (z/c)^{-\ell-1}$ (2.20)
with respect of the formal spectral parameter $u$.
Let ${\sf h}_{i}$, $i=1,\ldots,n$ be the operators
$\chi_{i}\ {\sf h}_{i}={\sf T}_{i,i}-\lambda_{i}[0]$ (2.21)
such that the vacuum vectors have zero eigenvalue
${\sf h}_{i}\ |0\rangle=0,\qquad\langle 0|\ {\sf h}_{i}=0\,.$ (2.22)
According to (2.11) and (2.12) the monodromy matrix entries $T_{k,l}(z)$ are eigenvector for the adjoint action of the operators ${\sf h}_{i}$
$[{\sf h}_{i},T_{k,l}(z)]=(\delta_{i,l}-\delta_{i,k})\ T_{k,l}(z)$ (2.23)
for $\mathfrak{gl}_{N}$-invariant monodromies and
$[{\sf h}_{i},T_{k,l}(z)]=(\delta_{i,l}-\delta_{i,k}+\delta_{i,-k}-\delta_{i,-l})\,T_{k,l}(z)$ (2.24)
for $\mathfrak{o}_{2n+1}$-invariant monodromies.
The Bethe vectors in the integrable model depends on a collection of sets
${\bar{t}}=\begin{cases}\{{\bar{t}}^{1},\ldots,{\bar{t}}^{n-1}\}\quad&\mbox{for}\quad\mathfrak{g}=\mathfrak{gl}_{n},\\ \{{\bar{t}}^{0},{\bar{t}}^{1},\ldots,{\bar{t}}^{n-1}\}\quad&\mbox{for}\quad\mathfrak{g}=\mathfrak{o}_{2n+1}\end{cases}$ (2.25)
of Bethe parameters ${\bar{t}}^{s}=\{{\bar{t}}^{s}_{1},\ldots,{\bar{t}}^{s}_{r_{s}}\}$ with cardinality $|{\bar{t}}^{s}|=r_{s}$. The superscript on the (sets of) Bethe parameters denotes the color of the (set of) Bethe parameters. The colors are in correspondence with the simple roots of the algebra $\mathfrak{g}$. The Bethe vectors themselves are certain polynomials of the monodromy entries $T_{i,j}(u)$ for $i\leq j$ acting to the right vacuum vector $|0\rangle$
${\mathbb{B}}(\bar{t})=\mathcal{P}\big{(}T_{i\leq j}(\bar{t})\big{)}\ |0\rangle=\mathcal{B}(\bar{t})\ |0\rangle,$ (2.26)
where the polynomial $\mathcal{B}(\bar{t})=\mathcal{P}\big{(}T_{i\leq j}(\bar{t})\big{)}$ is called a pre-Bethe vector.
Analogously, left or dual Bethe vectors are polynomials of monodromy entries $T_{i,j}(u)$ for $i\geq j$ acting to the left vacuum vector $|0\rangle$
${\mathbb{C}}(\bar{t})=\langle 0|\ \mathcal{P}(T_{i\geq j}(\bar{t}))=\langle 0|\ \mathcal{C}(\bar{t}).$ (2.27)
The structure of these polynomials can be fixed in the framework of the nested Bethe ansatz [15, 16] or by the method of projections [8, 9]. When the Bethe parameters are generic we call such Bethe vectors (2.26) and (2.27) off-shell Bethe vectors.
The off-shell Bethe vectors ${\mathbb{B}}({\bar{t}})$ and ${\mathbb{C}}({\bar{t}})$ are eigenvectors of the operators ${\sf h}_{i}$ with eigenvalues
${\sf h}_{i}\cdot{\mathbb{B}}({\bar{t}})=(r_{i-1}-r_{i})\ {\mathbb{B}}({\bar{t}}),\qquad{\mathbb{C}}({\bar{t}})\cdot{\sf h}_{i}=(r_{i-1}-r_{i})\ {\mathbb{C}}({\bar{t}}),\qquad i=1,\ldots,n.$ (2.28)
Remark 2.1.
In what follows, we will consider only the Bethe vectors ${\mathbb{B}}({\bar{t}})$. All the relations for the dual Bethe vectors ${\mathbb{C}}({\bar{t}})$ can be obtained from the corresponding relations for ${\mathbb{B}}({\bar{t}})$ using the transposition antihomomorphism. When we consider different embeddings we will use the notation $\mathbb{B}_{\mathfrak{gl}}({\bar{t}})$ and $\mathbb{B}_{\mathfrak{o}}({\bar{t}})$ to distinguish the off-shell Bethe vectors in the models with different symmetries. But most often we will the use notation ${\mathbb{B}}({\bar{t}})$ since it will be clear from the context what type of Bethe vector we are exploring.
Remark 2.2.
The commutation relations (2.5) between the monodromy matrix entries do not depend on the parameters $\chi_{i}$ explicitly, nor do the definition of the vacuum state (2.26). Since the Bethe vectors are polynomials in the entries of the monodromy matrix, they do not depend explicitly on the parameters $\chi_{i}$ either. It is only when using the expansion (2.6) that the dependence in the $\chi_{i}$ parameters becomes explicit, as for instance in the relations (2.11) or (2.12) which involve the zero mode action. Therefore, any relation involving only the Bethe vectors, the entries $T_{ij}(z)$ and/or the eigenvalues $\lambda_{i}(z)$ should not depend on the $\chi$’s. In other words, if the $\chi_{i}$ parameters appear explicitly in such relation, each coefficient of the parameters $\chi_{i}$ should be set to zero independently. We will use this property in the appendix B.
On-shell Bethe vectors
Due to the $RTT$ commutation relations (2.4) the trace of the monodromy matrix (the transfer matrix)
$\mathcal{T}(z)=\sum_{i\in I_{\mathfrak{g}}}T_{i,i}(z)$ (2.29)
commutes
$\mathcal{T}(u)\,\mathcal{T}(v)=\mathcal{T}(v)\,\mathcal{T}(u)$
at different values of the spectral parameters and generates a set of commuting operators. The Bethe vectors become eigenvectors of the transfer matrix (also known as on-shell Bethe vectors) if the Bethe parameters satisfy so called Bethe equations (see below (2.34). To describe the Bethe equations in the $\mathfrak{g}$-invariant integrable models we introduce the rational functions
$f(u,v)=1+g(u,v)=h(u,v)g(u,v)=\frac{u-v+c}{u-v},\qquad\mathfrak{f}(u,v)=\frac{u-v+c/2}{u-v}\,.$ (2.30)
We also define the functions
$\alpha_{s}(z)=\frac{\lambda_{s}(z)}{\lambda_{s+1}(z)},\qquad s=0,\ldots,n-1.$ (2.31)
We will use the following convention for the products of scalar functions depending on sets of parameters, for example
$\lambda_{s}({\bar{t}}^{s})=\prod_{t^{s}_{a}\in{\bar{t}}^{s}}\lambda_{s}(t^{s}_{a}),\qquad f({\bar{t}}^{s},{\bar{t}}^{s^{\prime}})=\prod_{t^{s}_{a}\in{\bar{t}}^{s}}\prod_{t^{s^{\prime}}_{b}\in{\bar{t}}^{s^{\prime}}}f(t^{s}_{a},t^{s^{\prime}}_{b}),\quad\mbox{etc.}$. (2.32)
with $f(\varnothing,{\bar{t}})=f({\bar{t}},\varnothing)=1$.
The on-shell Bethe vectors are eigenvectors of the transfer matrix
$\mathcal{T}(z)\cdot{\mathbb{B}}({\bar{t}})=\tau(z;{\bar{t}})\,{\mathbb{B}}({\bar{t}})$ (2.33)
if each set ${\bar{t}}^{s}$ of Bethe parameters satisfy the equalities
$\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I})=\frac{\lambda_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I})}{\lambda_{s+1}({\bar{t}}^{s}_{\scriptscriptstyle\rm I})}=\frac{f_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II})}{f_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm I})}\ \frac{f({\bar{t}}^{s+1},{\bar{t}}^{s}_{\scriptscriptstyle\rm I})}{f({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s-1})}\quad\text{with}\begin{cases}s=1,\ldots,n-1,\quad\text{for }\ \mathfrak{gl}_{n},\\ s=0,1,\ldots,n-1,\quad\text{for }\ \mathfrak{o}_{2n+1},\end{cases}$ (2.34)
for any disjoint partition ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}$ and with the conditions ${\bar{t}}^{0}={\bar{t}}^{n}=\varnothing$ for $\mathfrak{gl}_{n}$ and ${\bar{t}}^{-1}={\bar{t}}^{n}=\varnothing$ for $\mathfrak{o}_{2n+1}$. In (2.34), we have used the convention (2.32) for products of functions and the functions $f_{s}(u,v)$ are defined as
$f_{s}(u,v)$ $=$ $\begin{cases}\mathfrak{f}(u,v),\quad&s=0,\\ f(u,v),\quad&s=1,\ldots,n-1\,.\end{cases}$ (2.35)
For the $\mathfrak{gl}_{n}$-invariant models the eigenvalue $\tau(z;{\bar{t}})$ in (2.33) is equal to
$\tau(z;{\bar{t}})=\sum_{s=1}^{n}\lambda_{s}(z)\ f({\bar{t}}^{s},z)\ f(z,{\bar{t}}^{s-1})\,.$ (2.36)
In the case of $\mathfrak{o}_{2n+1}$-invariant models this eigenvalue is
$\tau(z;{\bar{t}})=\lambda_{0}(z)\ f({\bar{t}}^{0},z_{0})\ f(z,{\bar{t}}^{0})+\sum_{s=1}^{n}\Big{(}\lambda_{s}(z)\ f({\bar{t}}^{s},z)\ f(z,{\bar{t}}^{s-1})+\lambda_{-s}(z)\ f({\bar{t}}^{s-1},z_{s-1})\ f(z_{s},{\bar{t}}^{s})\Big{)}\,,$
where we have used the notations (2.19) and $\lambda_{-s}(z)$ satisfy the relations (2.18).
The recurrence relations for the off-shell Bethe vectors will be written as sums over partitions of the sets of Bethe parameters. A partition ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III}\}$ corresponds to a decomposition into (possibly empty) disjoint subsets ${\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III}$ such that ${\bar{t}}^{s}={\bar{t}}^{s}_{\scriptscriptstyle\rm I}\cup{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\cup{\bar{t}}^{s}_{\scriptscriptstyle\rm III}$ and ${\bar{t}}^{s}_{\scriptscriptstyle\rm I}\cap{\bar{t}}^{s}_{\scriptscriptstyle\rm II}={\bar{t}}^{s}_{\scriptscriptstyle\rm I}\cap{\bar{t}}^{s}_{\scriptscriptstyle\rm III}={\bar{t}}^{s}_{\scriptscriptstyle\rm II}\cap{\bar{t}}^{s}_{\scriptscriptstyle\rm III}=\varnothing$. The cardinalities of the subsets satisfy the equality
$|{\bar{t}}^{s}|=|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|+|{\bar{t}}^{s}_{\scriptscriptstyle\rm II}|+|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|\,,$
where some of the cardinalities $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|$, $|{\bar{t}}^{s}_{\scriptscriptstyle\rm II}|$, $|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|$ can be zero.
3 Bethe vectors in $\mathfrak{gl}_{n}$-invariant models
To denote a partial collection of sets of Bethe parameters $\{{\bar{t}}^{i},{\bar{t}}^{i+1},\ldots,{\bar{t}}^{j}\}$ for $i\leq j$ we will use the notation $\{{\bar{t}}^{s}\}_{i}^{j}$. For example, ${\bar{t}}=\{{\bar{t}}^{s}\}_{1}^{n-1}$. We will always assume that the partial collection of sets $\{{\bar{t}}^{s}\}_{i}^{j}$ is empty if $j<i$.
For $11\leq\ell<n-$ 1 and $1<k\leq n$ we introduce the functions
$\psi_{\ell}(z;{\bar{t}})=g(z,{\bar{t}}^{\ell-1})\,h({\bar{t}}^{\ell},z),\qquad\phi_{k}(z;{\bar{t}})=h(z,{\bar{t}}^{k-1})\,g({\bar{t}}^{k},z)\,.$ (3.1)
Remark that $\psi_{1}(z;{\bar{t}})=h({\bar{t}}^{1},z)$ and $\phi_{n}(z;{\bar{t}})=h(z,{\bar{t}}^{n-1})$ because $g(z,{\bar{t}}^{0})=g({\bar{t}}^{n},z)\equiv 1$ since ${\bar{t}}^{0}={\bar{t}}^{n}\equiv\varnothing$.
3.1 Action formula for $T_{1,n}(z)$ and ${\sf T}_{\ell+1,\ell}$
The off-shell Bethe vectors ${\mathbb{B}}({\bar{t}})$ in $\mathfrak{gl}_{n}$-invariant models are normalized in such a way that the action of the monodromy entry $T_{1,n}(z)$ onto off-shell Bethe vector ${\mathbb{B}}({\bar{t}})$ has the simplest form [^2] Note that normalization of the off-shell Bethe vectors used in (3.2) differs from the normalization used in the paper [7], but it is more suitable for the propose of this paper.
$T_{1,n}(z)\cdot{\mathbb{B}}({\bar{t}})=\mu_{1}^{n}(z;{\bar{t}})\,{\mathbb{B}}(\bar{u})\,,$ (3.2)
where $\bar{u}=\{\bar{u}^{1},\bar{u}^{2},\ldots,\bar{u}^{n-1}\}$ is an extended sets of Bethe parameters such that $\bar{u}^{s}=\{{\bar{t}}^{s},z\}$ for $s=1,\ldots,n-1$ and the normalization factor $\mu_{\ell}^{k}(z;{\bar{t}})$ is defined as follows for any $1\leq\ell<k\leq n$
$\mu_{\ell}^{k}(z;{\bar{t}})=\lambda_{k}(z)\,\psi_{\ell}(z;{\bar{t}})\,\phi_{k}(z;{\bar{t}})=\lambda_{k}(z)\,g(z,{\bar{t}}^{\ell-1})\,h({\bar{t}}^{\ell},z)\,h(z,{\bar{t}}^{k-1})\,g({\bar{t}}^{k},z)\,.$ (3.3)
We denote by the symbol $\mathcal{Z}_{\ell}^{k}$ the operation of adding a parameter $z$ to the sets ${\bar{t}}^{\ell},\ldots,{\bar{t}}^{k-1}$ of Bethe parameters in the off-shell Bethe vectors ${\mathbb{B}}({\bar{t}})$:
$\mathcal{Z}_{\ell}^{k}\cdot{\mathbb{B}}({\bar{t}})={\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},\{{\bar{t}}^{s},z\}_{\ell}^{k-1},\{{\bar{t}}^{s}\}_{k}^{n-1})\,.$ (3.4)
For example, using this notation the action (3.2) can be written as follows
$\mathcal{Z}_{1}^{n}\cdot{\mathbb{B}}({\bar{t}})=\frac{1}{\mu_{1}^{n}(z;{\bar{t}})}\,T_{1,n}(z)\cdot{\mathbb{B}}({\bar{t}})\,.$ (3.5)
To find the rectangular recurrence relations for the off-shell Bethe vectors ${\mathbb{B}}({\bar{t}})$ we will use the commutation relations (2.11) in the particular case
$[{\sf T}_{\ell+1,\ell},T_{i,j}(z)]=\chi_{\ell+1}\ \delta_{\ell+1,j}\ T_{i,\ell}(z)-\chi_{\ell}\ \delta_{\ell,i}\ T_{\ell+1,j}(z)$ (3.6)
and the action of the zero mode operators ${\sf T}_{\ell+1,\ell}$ onto off-shell Bethe vectors. These actions as well as the action (3.2) can be calculated in the framework of the projection method (see details in [9]). We will write the action of the zero modes in the form
$\begin{split}{\sf T}_{\ell+1,\ell}\cdot{\mathbb{B}}({\bar{t}})&=\sum_{\rm part}\Big{(}\chi_{\ell+1}\ \alpha_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\ \Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})-\\ &\quad-\chi_{\ell}\ \Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})\Big{)}\ {\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},\{{\bar{t}}^{s}\}_{\ell+1}^{n-1})\,,\end{split}$ (3.7)
where the sum in (3.7) goes over partitions ${\bar{t}}^{\ell}\vdash\{{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\}$ such that $|{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|=1$. The functions $\Omega^{L}$ and $\Omega^{R}$ in (3.7) are defined as follows
$\Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})=\gamma({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\ \frac{h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1})}{g({\bar{t}}^{\ell+1},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})},\qquad\Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})=\gamma({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II})\ \frac{h({\bar{t}}^{\ell+1},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1})}\,,$ (3.8)
where the function $\gamma(u,v)$ is
$\gamma(u,v)=\frac{f(u,v)}{h(u,v)h(v,u)}=\frac{g(u,v)}{h(v,u)}\,.$
3.2 Rectangular recurrence relations for $\mathfrak{gl}_{n}$
For $m\in{\mathbb{Z}}$ we define the step function $\Theta(m)$
$\Theta(m)=\begin{cases}1,\quad m\geq 0,\\ 0,\quad m<0.\end{cases}$ (3.9)
Proposition 3.1.
For any pair of positive integers $1\leq\ell<k\leq n$, the off-shell Bethe vector $\mathcal{Z}_{\ell}^{k}\cdot{\mathbb{B}}({\bar{t}})$ satisfies the rectangular recurrence relation
$\begin{split}\mathcal{Z}_{\ell}^{k}\cdot{\mathbb{B}}({\bar{t}})={\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},\{\bar{u}^{s}\}_{\ell}^{k-1},\{{\bar{t}}^{s}\}_{k}^{n-1})=\frac{1}{\mu^{k}_{\ell}(z;{\bar{t}})}\ \sum_{i=1}^{\ell}\sum_{j=k}^{n}\sum_{\rm part}\Xi^{\ell,k}_{i,j}(z;{\bar{t}})\ T_{i,j}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})\,,\end{split}$ (3.10)
where the functions $\Xi^{\ell,k}_{i,j}(z;{\bar{t}})$
$\Xi^{\ell,k}_{i,j}(z;{\bar{t}})=\psi_{\ell}(z;{\bar{t}}_{\scriptscriptstyle\rm I})\,\phi_{k}(z;{\bar{t}}_{\scriptscriptstyle\rm III})\prod_{s=i}^{\ell-1}\Omega^{R}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})\prod_{s=k}^{j-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm III})\ \Omega^{L}({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})$ (3.11)
depend on the partitions and the sum in (3.10) goes over partitions ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III}\}$ with cardinalities
$|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=\begin{cases}\Theta(s-i),&s<\ell,\\ 0,&s\geq\ell,\end{cases}\qquad|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=\begin{cases}0,&s\leq k-1,\\ \Theta(j-s-1),&s>k-1.\end{cases}$ (3.12)
The sets ${\bar{t}}^{\ell},\ldots,{\bar{t}}^{k-1}$ are not parted and in (3.10) ${\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\equiv{\bar{t}}^{\ell}$ and ${\bar{t}}^{k-1}_{\scriptscriptstyle\rm II}\equiv{\bar{t}}^{k-1}$. If the set ${\bar{t}}^{s}=\varnothing$ and according to (3.12) one may have $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=1$ for some $i$ or $|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=1$ for some $j$, then the terms in the r.h.s. of (3.10) for such values of the indices $i$ and $j$ have to be discarded.
Proof.
The proposition 3.1 will be proved in appendix A. $\square$
Remark that since $\ell<k$, the cardinalities in (3.12) show that for any color $s$, either ${\bar{t}}^{s}_{\scriptscriptstyle\rm I}$ or ${\bar{t}}^{s}_{\scriptscriptstyle\rm III}$ is empty. In other words, we always have partitions into a maximum of two subsets, ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}$ or ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III}\}$ with cardinalities $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|\leq 1$ and $|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|\leq 1$.
We call the recurrence relation (3.10) a rectangular recurrence relation because it implies the monodromy matrix entries $T_{i,j}(z)$ with $1\leq i\leq l$ and $k\leq j\leq n$. This defines a rectangular domain in monodromy matrix with vertices $T_{1,n}(z)$, $T_{l,n}(z)$, $T_{l,k}$ and $T_{1,k}(z)$. It lies above the diagonal of the monodromy matrix because $l<k$.
Corollary 3.1.
Considering the particular case $k=\ell+1$ in the recurrence relation (3.10), we get
$\begin{split}{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},\{{\bar{t}}^{\ell},z\},\{{\bar{t}}^{s}\}_{\ell+1}^{n-1})&=\sum_{i=1}^{\ell}\sum_{j=\ell+1}^{n}\sum_{\rm part}\frac{T^{+}_{i,j}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell-1},{\bar{t}}^{\ell},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{\ell+1}^{j-1},\{{\bar{t}}^{s}\}_{j}^{n-1})}{\lambda_{\ell+1}(z)\,g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II})h(z,{\bar{t}}^{\ell})h({\bar{t}}^{\ell},z)g({\bar{t}}^{\ell+1}_{\scriptscriptstyle\rm II},z)}\times\\ &\quad\times\prod_{s=i}^{\ell-1}\gamma({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II})\frac{h({\bar{t}}^{s+1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II})}\prod_{s=\ell+1}^{j-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm III})\gamma({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III})\frac{h({\bar{t}}^{s}_{\scriptscriptstyle\rm III},\bar{t}^{s-1}_{\scriptscriptstyle\rm II})}{g({\bar{t}}^{s+1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III})}\,.\end{split}$
Such type of recurrence relations for arbitrary $\ell$ was written for the first time in [19] for the case of integrable models associated to $U_{q}(\mathfrak{gl}_{n})$. In the case of $\mathfrak{gl}_{n}$-invariant integrable models these recurrence relations for the cases $\ell=1,k=2$ and $\ell=n-1,k=n$ were presented in [4].
3.3 Embedding $Y(\mathfrak{gl}_{a})\otimes Y(\mathfrak{gl}_{n-a})\hookrightarrow Y(\mathfrak{gl}_{n})$
Let us consider the rectangular recurrence relations (3.10) when the set of Bethe parameters ${\bar{t}}^{a}$ is empty for some given $a$ in $\{1,\ldots,n-1\}$. These reduced recurrence relations will imply the following proposition.
Proposition 3.2.
The off-shell Bethe vectors $\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})$ factorizes into the product of $\mathfrak{gl}_{a}$-type pre-Bethe vectors $\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1})$ and $\mathfrak{gl}_{n-a}$-type pre-Bethe vectors $\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})$ acting on the vacuum vector $|0\rangle$:
$\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})=\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1})\cdot\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\,|0\rangle\,.$ (3.13)
Proof.
Let us consider the recurrence relations (3.10) for two different ranges of the indices $\ell$ and $k$: $1\leq\ell<k\leq a$ and $a+1\leq\ell<k\leq n$. These ranges for $k$ and $\ell$ ensure that the set ${\bar{t}}^{a}$ remains empty. Since $\mathbb{B}_{\mathfrak{gl}}(\varnothing)=|0\rangle$ (i.e. $\mathcal{B}_{\mathfrak{gl}}(\varnothing)=1$) the recurrence relations will prove the statement of (3.13) by induction on the cardinalities.
  • $1\leq\ell<k\leq a$. The cardinality of the empty subset ${\bar{t}}^{a}_{\scriptscriptstyle\rm III}$ given by (3.12) is equal to $\Theta(j-a-1)$ which restricts the summation over $j$ to $k\leq j\leq a$. This means that the recurrence relation for $\mathcal{Z}^{k}_{\ell}\cdot\mathbb{B}_{\mathfrak{gl}}({\bar{t}})|_{{\bar{t}}^{a}=\varnothing}$ for $1\leq\ell<k\leq a$ implies only the monodromy entries $T_{i,j}(z)$ for $1\leq i\leq\ell$ and $k\leq j\leq a$. The recurrence relation $\mathcal{Z}^{a}_{1}\cdot\mathbb{B}_{\mathfrak{gl}}({\bar{t}})|_{{\bar{t}}^{a}=\varnothing}$ will have only one term (there is no summation over partitions) corresponding to the action of monodromy entry $T_{1,a}(z)$ on the off-shell Bethe vector. Then the action of the zero mode operators ${\sf T}_{\ell+1,\ell}$ for $\ell=1,\ldots,a-1$ onto Bethe vector (3.13)
    $\begin{split}{\sf T}_{\ell+1,\ell}\cdot\mathbb{B}_{\mathfrak{gl}}&(\{{\bar{t}}^{s}\}_{1}^{a-1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})=\sum_{\rm part}\Big{(}\chi_{\ell+1}\ \alpha_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\,\Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})\\ &-\chi_{\ell}\ \Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})\Big{)}\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},\{{\bar{t}}^{s}\}_{\ell+1}^{a-1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})\end{split}$
    does not affect the sets of Bethe parameters $\{{\bar{t}}^{s}\}_{a+1}^{n-1}$ and so coincides with the action of the zero mode operators onto the Bethe vectors in $\mathfrak{gl}_{a}$-invariant model. This shows that the recurrence relations for $\mathcal{Z}^{k}_{\ell}\cdot\mathbb{B}_{\mathfrak{gl}}({\bar{t}})|_{{\bar{t}}^{a}=\varnothing}$ for $1\leq\ell<k\leq a$ coincides with the $\mathfrak{gl}_{a}$ recurrence relation for the vector $\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}\}_{1}^{a-1})$.
  • $a+1\leq\ell<k\leq n$. In this case the cardinality of the empty subset ${\bar{t}}^{a}_{\scriptscriptstyle\rm I}=\varnothing$, given by the equalities (3.12), yields the restriction for the index $i\geq a+1$. The summation over $i$ and $j$ in the r.h.s. of the recurrence relation for $\mathcal{Z}^{k}_{\ell}\cdot\mathbb{B}_{\mathfrak{gl}}({\bar{t}})|_{{\bar{t}}^{a}=\varnothing}$ will be restricted to the range $a+1\leq i\leq\ell<k\leq j\leq n$. The Bethe vector $\mathcal{Z}^{n}_{a+1}\cdot\mathbb{B}_{\mathfrak{gl}}({\bar{t}})$ will be proportional to the action of the monodromy matrix entry $T_{a+1,n}(z)\cdot\mathbb{B}_{\mathfrak{gl}}({\bar{t}})$ (there is no sum over partitions in this action) and the action of the zero mode operators ${\sf T}_{\ell+1,\ell}$ for $\ell=a+1,\ldots,n-1$ onto the Bethe vector (3.13) coincides with $\mathfrak{gl}_{n-a}$-type actions (3.7). The recurrence relation (3.10) in this case will not affect the sets of Bethe parameters $\{{\bar{t}}^{s}\}_{1}^{a-1}$ and will coincide with the $\mathfrak{gl}_{n-a}$-type recurrence relation for the Bethe vector $\mathcal{Z}^{k}_{\ell}\cdot\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})$, see (3.10).
$\square$
Remark 3.1.
The proposition 3.2 is a direct consequence of the projection method, when the off-shell Bethe vectors are expressed in terms of the Cartan-Weyl generators of the quantum affine algebra $U_{q}(\mathfrak{gl}_{n})$ [8] or Yangian double $\mathcal{D}Y(\mathfrak{gl}_{n})$ [9].
Splitting property of the Bethe vectors. Let us define the operation $\mathcal{Z}^{a+1}_{a}(t^{a})$ as the operation which adds the parameter $t^{a}$ to the set ${\bar{t}}^{a}$ of the Bethe vector $\mathbb{B}_{\mathfrak{gl}}({\bar{t}})$ according to definition (3.4). A direct consequence of the proposition 3.2 and the rectangular recurrence relations (3.10) is the presentation of the Bethe vector $\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}^{a-1}_{1},t^{a},\{{\bar{t}}^{s}\}_{a+1}^{n-1})$ in the form
$\begin{split}&\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}^{a-1}_{1},t^{a},\{{\bar{t}}^{s}\}_{a+1}^{n-1})=\mathcal{Z}^{a+1}_{a}(t^{a})\Big{(}\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1})\cdot\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\Big{)}\,|0\rangle=\\ &\quad=\frac{1}{\mu^{k}_{\ell}(t^{a};{\bar{t}})}\ \sum_{i=1}^{a}\sum_{j=a+1}^{n}\sum_{\rm part}\Xi^{\ell,k}_{i,j}(t^{a};{\bar{t}})\ T_{i,j}(t^{a})\cdot\Big{(}\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1})\cdot\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\Big{)}\,|0\rangle\,.\end{split}$ (3.14)
Iterating the relation (3.14) one gets the following presentation for the off-shell Bethe vector in the generic $\mathfrak{gl}_{n}$-invariant integrable model
$\mathbb{B}_{\mathfrak{gl}}({\bar{t}})=\mathcal{Z}^{a+1}_{a}(t^{a}_{1})\cdot\mathcal{Z}^{a+1}_{a}(t^{a}_{2})\cdots\mathcal{Z}^{a+1}_{a}(t^{a}_{r_{a}})\Big{(}\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{1}^{a-1})\cdot\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\Big{)}\,|0\rangle\,.$ (3.15)
The relation (3.15) was called in the papers [13, 14] the splitting property of the off-shell Bethe vectors, see e.g. proposition 3.1 in [14]. One can also compare the relation (3.13) with the equation (3.9) in [14]. These formulas were proved in these papers by a different approach using the evaluation homomorphism for the Yangian $Y(\mathfrak{gl}_{n})$ and the trace formula for the off-shell Bethe vectors.
4 Bethe vectors in $\mathfrak{o}_{2n+1}$-invariant models
The off-shell Bethe vectors in $\mathfrak{o}_{2n+1}$-invariant models depend on the sets of Bethe parameters ${\bar{t}}=({\bar{t}}^{0},{\bar{t}}^{1},\ldots,{\bar{t}}^{n-1})$ with cardinalities $|{\bar{t}}^{s}|=r_{s}$ for $s=0,1,\ldots,n-1$.
We normalize the $\mathfrak{o}_{2n+1}$-invariant off-shell Bethe vectors in such a way that the action of the monodromy matrix element $T_{-n,n}(z)$ onto ${\mathbb{B}}({\bar{t}})$ is given by the equality
$T_{-n,n}(z)\cdot{\mathbb{B}}({\bar{t}})=\mu_{-n}^{n}(z;{\bar{t}})\,{\mathbb{B}}(\bar{w})\,,$ (4.1)
where $\bar{w}=(\bar{w}^{0},\bar{w}^{1},\ldots,\bar{w}^{n-1})$ is a collection of extended sets of Bethe parameters such that
$\bar{w}^{s}=\{{\bar{t}}^{s},z,z_{s}\},\qquad z_{s}=z-c(s-1/2)$ (4.2)
and the normalization factor $\mu_{-n}^{n}(z;{\bar{t}})$ is given by the expression
$\mu_{-n}^{n}(z;{\bar{t}})=-\kappa_{n}\ \lambda_{n}(z)\ \frac{h(z,\bar{t}^{n-1})}{g(z_{n},\bar{t}^{n-1})}\ \frac{g(z_{1},\bar{t}^{0})}{h(z,\bar{t}^{0})}\,.$ (4.3)
Besides the sets $\bar{w}^{s}$ of cardinalities $r_{s}+2$ which are defined in (4.2) we will also use the sets $\bar{u}^{s}$ and $\bar{v}^{s}$ of cardinalities $r_{s}+1$ given by
$\bar{u}^{s}=\{{\bar{t}}^{s},z\},\qquad\bar{v}^{s}=\{{\bar{t}}^{s},z_{s}\},\qquad\bar{w}^{s}=\{\bar{u}^{s},z_{s}\}=\{\bar{v}^{s},z\}\,.$ (4.4)
For $-n\leq\ell<k\leq n$ let $\mathcal{Z}_{\ell}^{k}$ be an operator which extend the sets of Bethe parameters depending on the values of $\ell$ and $k$ according to the rules
$\mathcal{Z}_{\ell}^{k}\cdot{\mathbb{B}}({\bar{t}})=\begin{cases}{\mathbb{B}}(\{{\bar{t}}^{s}\}_{0}^{\ell-1},\{\bar{u}^{s}\}_{\ell}^{k-1},\{{\bar{t}}^{s}\}_{k}^{n-1}),&\qquad 0\leq\ell<k\leq n,\\[5.69054pt] {\mathbb{B}}(\{{\bar{t}}^{s}\}_{0}^{-k-1},\{\bar{v}^{s}\}_{-k}^{-\ell-1},\{{\bar{t}}^{s}\}_{-\ell}^{n-1}),&\qquad 0\leq-k<-\ell\leq n,\\[5.69054pt] {\mathbb{B}}(\{\bar{w}^{s}\}_{0}^{-\ell-1},\{\bar{u}^{s}\}_{-\ell}^{k-1},\{{\bar{t}}^{s}\}_{k}^{n-1}),&\qquad 0\leq-\ell\leq k\leq n,\\[5.69054pt] {\mathbb{B}}(\{\bar{w}^{s}\}_{0}^{k-1},\{\bar{v}^{s}\}_{k}^{-\ell-1},\{{\bar{t}}^{s}\}_{-\ell}^{n-1}),&\qquad 0\leq k\leq-\ell\leq n.\end{cases}$ (4.5)
The action (4.1) of the monodromy matrix entry $T_{-n,n}(z)$ can be presented as the action of the operator $\mathcal{Z}_{-n}^{n}$ onto off-shell Bethe vector
$\mathcal{Z}_{-n}^{n}\cdot{\mathbb{B}}({\bar{t}})=\frac{1}{\mu_{-n}^{n}(z;{\bar{t}})}\ T_{-n,n}(z)\cdot{\mathbb{B}}({\bar{t}})\,.$ (4.6)
The commutation relations between the zero mode operators ${\sf T}_{\ell+1,\ell}$ and the monodromy matrix elements take the form for $10\leq\ell\leq n-$ 1
$[{\sf T}_{\ell+1,\ell}\,,T_{i,j}(z)]=\chi_{\ell+1}\Big{(}\delta_{\ell,j-1}-\delta_{\ell,-j}\Big{)}\,T_{i,j-1}(z)-\chi_{\ell}\Big{(}\delta_{\ell,i}-\delta_{\ell,-i-1}\Big{)}\,T_{i+1,j}(z)\,.$ (4.7)
The action of the monodromy entry $T_{-n,n}(z)$ (4.1) as well as the action of the zero mode operators ${\sf T}_{\ell+1,\ell}$ onto off-shell Bethe vectors can be calculated in the framework of the projection method (see details in [18]). The action of the zero mode operators is
$\begin{split}{\sf T}_{\ell+1,\ell}\cdot{\mathbb{B}}({\bar{t}})=\sum_{\rm part}\Big{(}&\chi_{\ell+1}\ \alpha_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\ \Omega^{L}_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})\\ &-\chi_{\ell}\ \Omega^{R}_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})\Big{)}\ {\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},\{{\bar{t}}^{s}\}_{\ell+1}^{n-1})\,,\end{split}$ (4.8)
where the sum is over partitions with cardinalities $|{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|=1$. In (4.8) the rational functions $\Omega^{L}_{\ell}$ and $\Omega^{R}_{\ell}$ have expressions similar to (3.8):
$\begin{split}\Omega^{L}_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})&=\gamma_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\ \frac{h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1})}{g({\bar{t}}^{\ell+1},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})},\\ \Omega^{R}_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})&=\gamma_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II})\ \frac{h({\bar{t}}^{\ell+1},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1})}\,,\end{split}$ (4.9)
where
$\gamma_{\ell}(t_{1},t_{2})=\begin{cases}\displaystyle\mathfrak{f}(t_{1},t_{2})=\frac{t_{1}-t_{2}+c/2}{t_{1}-t_{2}},&\ell=0,\\[8.61108pt] \displaystyle\gamma(t_{1},t_{2})=\frac{g(t_{1},t_{2})}{h(t_{2},t_{1})}=\frac{c^{2}}{(t_{1}-t_{2})(t_{2}-t_{1}+c)},\quad&\ell=1,\ldots,n-1.\end{cases}$ (4.10)
4.1 Rectangular recurrence relations for $\mathfrak{o}_{2n+1}$
The main result of the paper is formulated in the theorem 4.1 below. The proof of this theorem is similar to the proof of the proposition (3.1) given in the appendix A. In the appendix B we will sketch its proof.
For the formulation of the statement of the theorem we need following notations:
  • the sign factor $\sigma_{m}$
    $\sigma_{m}=2\Theta(m-1)-1=\begin{cases}1,\quad m>0,\\ -1,\quad m\leq 0\end{cases}$ (4.11)
    which satisfy the property
    $\sigma_{m+1}=-\sigma_{-m}\,,$ (4.12)
  • the functions $\psi_{\ell}(z;{\bar{t}})$, $\phi_{k}(z;{\bar{t}})$, $\mu^{k}_{\ell}(z;{\bar{t}})$
    $\psi_{\ell}(z;{\bar{t}})=\begin{cases}g(z,{\bar{t}}^{\ell-1})\ h({\bar{t}}^{\ell},z),&0<\ell<n,\\[1.80832pt] g(z_{0},{\bar{t}}^{0}),&\ell=0,\\ \displaystyle{\frac{g({\bar{t}}^{-\ell},z_{-\ell})}{g(z_{-\ell},{\bar{t}}^{-\ell-1})}},&-n\leq\ell<0,\end{cases}$ (4.13)
    $\phi_{k}(z;{\bar{t}})=\begin{cases}h(z,{\bar{t}}^{k-1})\ g({\bar{t}}^{k},z),&0<k\leq n,\\[1.80832pt] g(z,{\bar{t}}^{0}),&k=0,\\ \displaystyle{\frac{g(z_{-k-1},{\bar{t}}^{-k-1})}{g({\bar{t}}^{-k},z_{-k-1})}},&-n<k<0,\end{cases}$ (4.14)
    $\mu^{k}_{\ell}(z;{\bar{t}})=\sigma_{-\ell-k}\ (\kappa_{k})^{\delta_{k,-\ell}}\ \lambda_{k}(z)\ \psi_{\ell}(z;{\bar{t}})\ \phi_{k}(z;{\bar{t}})\ \Big{(}\frac{g(z_{1},{\bar{t}}^{0})}{h(z,{\bar{t}}^{0})}\Big{)}^{\delta_{\ell<0,\,k>0}}\,,$ (4.15)
    where $\kappa_{k}=k-1/2$ and $\delta_{\rm condition}$ is equal to 1 if ”condition” is satisfied and to 0 otherwise.
  • the partitions
    $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=\begin{cases}\Theta(\ell)\Big{(}\Theta(s-i)+\Theta(-i-s-1)\Big{)},\quad s<|\ell|,\\ \Theta(-i-s-1),\quad s\geq|\ell|,\end{cases}$ (4.16)
    $|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=\begin{cases}\Theta(-k)\Big{(}\Theta(j+s)+\Theta(j-s-1)\Big{)},\quad s<|k|,\\ \Theta(j-s-1),\quad s\geq|k|,\end{cases}$ (4.17)
    $|{\bar{t}}^{s}_{\scriptscriptstyle\rm II}|=|{\bar{t}}^{s}|-|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|-|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|\,,$ (4.18)
  • the functions $\Gamma^{L}_{i,j}({\bar{t}}_{\scriptscriptstyle\rm II},{\bar{t}}_{\scriptscriptstyle\rm I})$ and $\Gamma^{R}_{i,j}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{\scriptscriptstyle\rm II})$ which depend on the partitions
    $\begin{split}\Gamma^{L}_{i,j}({\bar{t}}_{\scriptscriptstyle\rm II},{\bar{t}}_{\scriptscriptstyle\rm I})&=\prod_{s=i}^{j-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I})\Omega^{L}_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II}),\\ \Gamma^{R}_{i,j}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{\scriptscriptstyle\rm II})&=\prod_{s=i}^{j-1}\Omega^{R}_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II}),\end{split}$ (4.19)
    and the functions $\Xi^{\ell,k}_{i,j}(z;{\bar{t}})$
    $\begin{split}\Xi^{\ell,k}_{i,j}(z;{\bar{t}})&=\psi_{\ell}(z,{\bar{t}}_{\scriptscriptstyle\rm I})\ \phi_{k}(z,{\bar{t}}_{\scriptscriptstyle\rm III})\ \Gamma^{R}_{0,n}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{{\scriptscriptstyle\rm II},{\scriptscriptstyle\rm III}})\ \Gamma^{L}_{0,n}({\bar{t}}_{\scriptscriptstyle\rm II},{\bar{t}}_{{\scriptscriptstyle\rm III}})=\\ &=\psi_{\ell}(z,{\bar{t}}_{\scriptscriptstyle\rm I})\ \phi_{k}(z,{\bar{t}}_{\scriptscriptstyle\rm III})\ \Gamma^{R}_{0,n}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{{\scriptscriptstyle\rm II}})\ \Gamma^{L}_{0,n}({\bar{t}}_{{\scriptscriptstyle\rm I},{\scriptscriptstyle\rm II}},{\bar{t}}_{{\scriptscriptstyle\rm III}})\,,\end{split}$ (4.20)
    where the equality between first and second lines in (4.20) directly follows from the definitions (4.9). Note that $\Gamma^{L}_{i,j}({\bar{t}}_{\scriptscriptstyle\rm II},{\bar{t}}_{\scriptscriptstyle\rm I})=\Gamma^{R}_{i,j}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{\scriptscriptstyle\rm II})\equiv 1$ for $i\geq j$.
Theorem 4.1.
For any $\ell,k$ such that $-n\leq l<k\leq n$ the off-shell Bethe vector $\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})$ satisfies the rectangular recurrence relation
$\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})=\frac{1}{\mu^{k}_{\ell}(z;{\bar{t}})}\sum_{i=-n}^{\ell}\sum_{j=k}^{n}\sum_{\rm part}\Xi^{\ell,k}_{i,j}(z;{\bar{t}})\ (\sigma_{-i})^{\delta_{\ell\geq 0}}(\sigma_{j})^{\delta_{k\leq 0}}\ T_{i,j}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})\,,$ (4.21)
where the sum is over partitions ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III}\}$ with cardinalities depending on the indices $-n\leq i\leq\ell<k\leq j\leq n$ and given by the equalities (4.16), (4.17) and (4.18). If the cardinality of the set ${\bar{t}}^{s}$ is small and according to (4.16) and (4.17), one may have $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|+|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|>|{\bar{t}}^{s}|$ for some $i$ and $j$. Then the terms in the r.h.s. of (4.21) for such values of the indices $i$ and $j$ have to be discarded.
Let us stress that, in opposition to the $\mathfrak{gl}_{n}$ case, for $\mathfrak{o}_{2n+1}$-invariant models, depending on the color $s$, the subsets ${\bar{t}}^{s}_{\scriptscriptstyle\rm I}$ and ${\bar{t}}^{s}_{\scriptscriptstyle\rm III}$ can be both non-empty at the same time, and the cardinalities of ${\bar{t}}^{s}_{\scriptscriptstyle\rm I}$ and ${\bar{t}}^{s}_{\scriptscriptstyle\rm III}$ can be up to 2.
4.2 Examples of recurrence relations
We provide subcases of the recurrence relations (4.21) that have some relevance for the study of integrable models.
Elementary recurrence relations
For explicit calculations, the most interesting (and simplest) cases of the recurrence relations (4.21) are the $2n$ cases when $k=\ell+1$ and $\ell=-n,\ldots,n-1$. These $2n$ cases are gathered in four classes of expressions, depending on the value of the index $\ell$.
  • When $0<\ell<n$, the recurrence relation reads
    $\begin{split}&\mathcal{Z}^{\ell+1}_{\ell}\cdot{\mathbb{B}}({\bar{t}})={\mathbb{B}}({\bar{t}}^{0},\ldots,{\bar{t}}^{\ell-1},\{{\bar{t}}^{\ell},z\},{\bar{t}}^{\ell+1},\ldots,{\bar{t}}^{n-1})=\\ &\quad=\sum_{i=-n}^{\ell}\sum_{j=\ell+1}^{n}\sum_{\rm part}\frac{g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\,h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)\,g({\bar{t}}^{\ell+1}_{\scriptscriptstyle\rm III},z)}{\lambda_{\ell+1}(z)\,g(z,{\bar{t}}^{\ell-1})\,h({\bar{t}}^{\ell},z)\,h(z,{\bar{t}}^{\ell})\,g({\bar{t}}^{\ell+1},z)}\times\\ &\qquad\times\Gamma^{R}_{[i],\ell_{i}}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{{\scriptscriptstyle\rm II}})\ \Gamma^{L}_{\ell+1,j}({\bar{t}}_{{\scriptscriptstyle\rm I},{\scriptscriptstyle\rm II}},{\bar{t}}_{{\scriptscriptstyle\rm III}})\ \sigma_{i+1}\ T_{i,j}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})\,,\end{split}$ (4.22)
    where $\ell_{i}={\rm max}(\ell,|i|)$, $[i]=\frac{1}{2}(i+|i|)$, and the cardinalities of the subsets
    $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=\begin{cases}\Theta(s-i)+\Theta(-i-s-1),&s<\ell,\\ \Theta(-i-s-1),&s\geq\ell,\end{cases}\qquad|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=\begin{cases}0,&s<\ell+1,\\ \Theta(j-s-1),&s\geq\ell+1.\end{cases}$ (4.23)
    Remark that for $\ell=n-1$ this recurrence relation reduces to
    $\begin{split}{\mathbb{B}}(\{t^{s}\}_{0}^{n-2},\{{\bar{t}}^{n-1},z\})&=\frac{1}{\lambda_{n}(z)\ h(z,{\bar{t}}^{n-1})}\times\\ &\times\sum_{i=-n}^{n-1}\sum_{\rm part}\frac{\sigma_{i+1}\,T_{i,n}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})}{g(z,{\bar{t}}^{n-2}_{\scriptscriptstyle\rm II})\,h({\bar{t}}^{n-1}_{\scriptscriptstyle\rm II},z)}\prod_{s=0}^{n-1}\gamma_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II})\prod_{s=1}^{n-1}\frac{h({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II})}\end{split}$
    which coincides with the recurrence relation (3.16) from the paper [18].
  • When $0<\ell<n$, we get a so called shifted recurrence relation for the off-shell Bethe vectors. It corresponds to an extension of the set of Bethe parameters ${\bar{t}}^{\ell}$ by the shifted parameter $z_{\ell}=z-c(\ell-1/2)$:
    $\begin{split}&\mathcal{Z}^{-\ell}_{-\ell-1}\cdot{\mathbb{B}}({\bar{t}})={\mathbb{B}}({\bar{t}}^{0},\ldots,{\bar{t}}^{\ell-1},\{{\bar{t}}^{\ell},z_{\ell}\},{\bar{t}}^{\ell+1},\ldots,{\bar{t}}^{n-1})=\\ &\quad=\sum_{i=-n}^{-\ell-1}\sum_{j=-\ell}^{n}\sum_{\rm part}\frac{(-1)^{\delta_{j\geq\ell+1}}\ g({\bar{t}}^{\ell+1}_{\scriptscriptstyle\rm I},z_{\ell+1})\,h(z_{\ell},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm III})\,g(z_{\ell-1},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm III})}{\lambda_{-\ell}(z)\,g({\bar{t}}^{\ell+1},z_{\ell+1})\,h({\bar{t}}^{\ell},z_{\ell})\,h(z_{\ell},{\bar{t}}^{\ell})\,g(z_{\ell-1},{\bar{t}}^{\ell-1})}\times\\ &\qquad\times\Gamma^{R}_{\ell+1,-i}({\bar{t}}_{\scriptscriptstyle\rm I},{\bar{t}}_{{\scriptscriptstyle\rm II},{\scriptscriptstyle\rm III}})\ \Gamma^{L}_{[-j],\ell_{j}}({\bar{t}}_{\scriptscriptstyle\rm II},{\bar{t}}_{{\scriptscriptstyle\rm III}})\ \sigma_{j}\ T_{i,j}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})\,,\end{split}$ (4.24)
    where $\ell_{j}={\rm max}(\ell,|j|)$, $[j]=\frac{1}{2}(j+|j|)$ as above, and the cardinalities of the subsets in the sum over partitions are
    $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=\begin{cases}0,&s<\ell+1,\\ \Theta(-i-s-1),&s\geq\ell+1,\end{cases}\qquad|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=\begin{cases}\Theta(s+j)+\Theta(j-s-1),&s<\ell,\\ \Theta(j-s-1),&s\geq\ell.\end{cases}$ (4.25)
    The sign factor $(-1)^{\delta_{j\geq\ell+1}}$ in the second line of (4.24) is due to the identity $g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm III},z_{\ell-1})^{-1}=-h(z_{\ell},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm III})$, valid when the cardinality of the subset ${\bar{t}}^{\ell}_{\scriptscriptstyle\rm III}$ is equal to one. According to (4.25) this happens when $j\geq\ell+1$.
    For $\ell=n-1$ the recurrence relation (4.24) can be rewritten in the form
    $\begin{split}&{\mathbb{B}}(\{{\bar{t}}^{s}\}_{0}^{n-2},\{{\bar{t}}^{n-1},z_{n-1}\})=\frac{1}{\lambda_{-n+1}(z)\ h({\bar{t}}^{n-1},z_{n-1})}\times\\ &\quad\times\sum_{j=-n+1}^{n}\sum_{\rm part}\frac{(-1)^{\delta_{j,n}}}{h(z_{n-1},{\bar{t}}^{n-1}_{\scriptscriptstyle\rm II})}\frac{\sigma_{j}\,T_{-n,j}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})}{g(z_{n-2},{\bar{t}}^{n-2}_{\scriptscriptstyle\rm II})}\prod_{s=0}^{n-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm III})\gamma_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III})\prod_{s=1}^{n-1}\frac{h({\bar{t}}^{s}_{\scriptscriptstyle\rm III},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II})}{g({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm III})}\end{split}$ (4.26)
    and coincides exactly with the recurrence relation (3.19) from the paper [18]. The sign factor $(-1)^{\delta_{j\geq\ell+1}}$ turns in this case to the sign factor $(-1)^{\delta_{j,n}}$.
  • When $\ell=0$, the recurrence relation takes the form
    $\begin{split}\mathcal{Z}^{1}_{0}\cdot{\mathbb{B}}({\bar{t}})&={\mathbb{B}}(\{{\bar{t}}^{0},z\},{\bar{t}}^{1},\ldots,{\bar{t}}^{n-1})=\sum_{i=-n}^{0}\sum_{j=1}^{n}\sum_{\rm part}\frac{g(z_{0},{\bar{t}}^{0}_{\scriptscriptstyle\rm I})\,g({\bar{t}}^{1}_{\scriptscriptstyle\rm III},z)\,\sigma_{i+1}\ T_{i,j}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})}{\lambda_{1}(z)\,g(z_{0},{\bar{t}}^{0})\,h(z,{\bar{t}}^{0})\,g({\bar{t}}^{1},z)}\times\\ &\qquad\times\prod_{s=0}^{-i-1}\gamma_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II})\frac{h({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II})}\prod_{s=1}^{j-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm III})\gamma_{s}({\bar{t}}^{s}_{{\scriptscriptstyle\rm I},{\scriptscriptstyle\rm II}},{\bar{t}}^{s}_{\scriptscriptstyle\rm III})\frac{h({\bar{t}}^{s}_{\scriptscriptstyle\rm III},{\bar{t}}^{s-1}_{{\scriptscriptstyle\rm I},{\scriptscriptstyle\rm II}})}{g({\bar{t}}^{s}_{{\scriptscriptstyle\rm I},{\scriptscriptstyle\rm II}},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm III})}\,,\end{split}$ (4.27)
    where cardinalities of the subsets in the sum over partitions are
    $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=\Theta(-i-s-1),\quad\forall s,\qquad|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=\begin{cases}0,\quad&s=0,\\ \Theta(j-s-1),\quad&s\geq 1.\end{cases}$ (4.28)
  • When $\ell=-1$ one gets the recurrence relation
    $\begin{split}\mathcal{Z}^{0}_{-1}\cdot{\mathbb{B}}({\bar{t}})&={\mathbb{B}}(\{{\bar{t}}^{0},z_{0}\},{\bar{t}}^{1},\ldots,{\bar{t}}^{n-1})=\sum_{i=-n}^{-1}\sum_{j=0}^{n}\sum_{\rm part}\frac{g({\bar{t}}^{1}_{\scriptscriptstyle\rm I},z_{1})\,g(z,{\bar{t}}^{0}_{\scriptscriptstyle\rm III})\ \sigma_{j}\,T_{i,j}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})}{\lambda_{0}(z)\,g({\bar{t}}^{1},z_{1})\,g(z_{1},{\bar{t}}^{0})^{-1}\,g(z,{\bar{t}}^{0})}\times\\ &\qquad\times\prod_{s=1}^{-i-1}\gamma_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{{\scriptscriptstyle\rm II},{\scriptscriptstyle\rm III}})\frac{h({\bar{t}}^{s}_{{\scriptscriptstyle\rm II},{\scriptscriptstyle\rm III}},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s-1}_{{\scriptscriptstyle\rm II},{\scriptscriptstyle\rm III}})}\prod_{s=0}^{j-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm III})\gamma_{s}({\bar{t}}^{s}_{{\scriptscriptstyle\rm II}},{\bar{t}}^{s}_{\scriptscriptstyle\rm III})\frac{h({\bar{t}}^{s}_{\scriptscriptstyle\rm III},{\bar{t}}^{s-1}_{{\scriptscriptstyle\rm II}})}{g({\bar{t}}^{s}_{{\scriptscriptstyle\rm II}},{\bar{t}}^{s-1}_{\scriptscriptstyle\rm III})}\,,\end{split}$ (4.29)
    where cardinalities of the subsets in the sum over partitions are
    $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=\begin{cases}0,&s=0,\\ \Theta(-i-s-1),\quad&s\geq 1,\end{cases}\qquad|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=\Theta(j-s-1),\quad\forall s\geq 1.$ (4.30)
Case of $\mathfrak{o}_{3}$-invariant models
The above recurrence relations were calculated for the algebras $\mathfrak{o}_{2n+1}$ with $n>1$. However, one can verify that they are also valid for the most simple orthogonal algebra $\mathfrak{o}_{3}$ when $n=1$. This case was investigated in details in the paper [20]. In this case there are only three cases of rectangular recurrence relations for the Bethe vectors: $\mathcal{Z}^{1}_{-1}\cdot{\mathbb{B}}({\bar{t}}^{0})$, $\mathcal{Z}^{1}_{0}\cdot{\mathbb{B}}({\bar{t}}^{0})$ and $\mathcal{Z}^{0}_{-1}\cdot{\mathbb{B}}({\bar{t}}^{0})$, that we detail below.
For the Bethe vector $\mathcal{Z}^{1}_{-1}\cdot{\mathbb{B}}({\bar{t}}^{0})$, there is no summation over partitions and according to (4.15) it is equal to
$\mathcal{Z}^{1}_{-1}\cdot{\mathbb{B}}({\bar{t}}^{0})={\mathbb{B}}(\{{\bar{t}}^{0},z,z_{0}\})=-\frac{\lambda_{1}(z)}{2}\ T_{-1,1}(z)\cdot{\mathbb{B}}({\bar{t}}^{0})\,.$ (4.31)
For the Bethe vector $\mathcal{Z}^{1}_{0}\cdot{\mathbb{B}}({\bar{t}}^{0})$ the recurrence relation (4.27) becomes
$\begin{split}&\mathcal{Z}^{1}_{0}\cdot{\mathbb{B}}({\bar{t}}^{0})={\mathbb{B}}(\{{\bar{t}}^{0},z\})=\\ &=\frac{1}{\lambda_{1}(z)\mathfrak{f}(z_{0},{\bar{t}}^{0})}\left(T_{0,1}(z)\cdot{\mathbb{B}}({\bar{t}}^{0})-\sum_{\rm part}g(z_{0},{\bar{t}}^{0}_{\scriptscriptstyle\rm I})\mathfrak{f}({\bar{t}}^{0}_{\scriptscriptstyle\rm I},{\bar{t}}^{0}_{\scriptscriptstyle\rm II})\ T_{-1,1}(z)\cdot{\mathbb{B}}({\bar{t}}^{0}_{\scriptscriptstyle\rm II})\right)\,,\end{split}$ (4.32)
while for the Bethe vector $\mathcal{Z}^{0}_{-1}\cdot{\mathbb{B}}({\bar{t}}^{0})$ the recurrence relation (4.29) becomes
$\begin{split}&\mathcal{Z}^{0}_{-1}\cdot{\mathbb{B}}({\bar{t}}^{0})={\mathbb{B}}(\{{\bar{t}}^{0},z_{0}\})=\\ &=\frac{1}{\lambda_{0}(z)\mathfrak{f}({\bar{t}}^{0},z)}\left(-T_{-1,0}(z)\cdot{\mathbb{B}}({\bar{t}}^{0})+\sum_{\rm part}g(z,{\bar{t}}^{0}_{\scriptscriptstyle\rm III})\,\alpha_{0}({\bar{t}}^{0}_{\scriptscriptstyle\rm III})\,\mathfrak{f}({\bar{t}}^{0}_{\scriptscriptstyle\rm II},{\bar{t}}^{0}_{\scriptscriptstyle\rm III})\ T_{-1,1}(z)\cdot{\mathbb{B}}({\bar{t}}^{0}_{\scriptscriptstyle\rm II})\right)\,.\end{split}$ (4.33)
In (4.32) and (4.33) we have $|{\bar{t}}^{0}_{\scriptscriptstyle\rm I}|=|{\bar{t}}^{0}_{\scriptscriptstyle\rm III}|=1$. Formulas (4.31), (4.32) and (4.33) coincide with the formulas (4.15), (4,19) and (4.20) of the paper [20]. It proves that the theorem 4.1 is also valid for $\mathfrak{o}_{3}$-invariant integrable models.
Remark 4.1.
The $n=1$ example shows that for $\mathfrak{o}_{2n+1}$-invariant integrable models, there is no simple symmetry transformation which relates the shifted (4.24) and non-shifted (4.22) recurrence relations. This is in opposition to the $\mathfrak{gl}_{n}$-integrable models, where there is a symmetry relation for Bethe vectors, see [21] and the $\mathfrak{gl}_{n}$ reduction in section 4.3.
Indeed, one can replace $z\to z_{0}$ in (4.32) so the l.h.s. of (4.32) and (4.33) coincide. But then, in order to prove that the r.h.s. of these equalities also coincide, one has to use the following relation between monodromy matrix entries
$\begin{split}T_{-1,0}(z)&=-T_{0,1}(z_{0})\,T_{1,1}(z_{0})^{-1}\,T_{0,0}(z)+\\ &\quad+\ \Big{(}T_{0,1}(z_{0})\,T_{1,1}(z_{0})^{-1}\,T_{0,1}(z)\,T_{1,1}(z)^{-1}+T_{-1,1}(z)\,T_{1,1}(z)^{-1}\Big{)}\,T_{1,0}(z)\,,\end{split}$ (4.34)
which is a direct consequence of (2.9). Then, in principle, using the action formulas for $T_{0,0}(z)$, $T_{1,1}(z)^{-1}$ and $T_{1,0}(z)$ one could connect the shifted and non-shifted recurrence formula, but these calculations are too complicated to be carried in the general case. Action formulas for $T_{0,0}(z)$, $T_{1,1}(z)$ and $T_{1,0}(z)$ have been computed in [21] for $n=1$ and in [18] for arbitrary $n$.
In conclusion, in $\mathfrak{o}_{2n+1}$-invariant quantum integrable models, there is no simple answer to the question of the existence of a symmetry transformation in the space of the off-shell Bethe vectors based on the symmetry relation (2.9) which relates the non-shifted and shifted recurrence relations.
4.3 Reduction to $\mathfrak{gl}_{n}$
Still in the framework of the $\mathfrak{o}_{2n+1}$-type monodromy matrix, we consider the particular case of Bethe vectors with ${\bar{t}}^{0}=\varnothing$. In that case, since the color 0 is not implied, one should recover the construction done for the $\mathfrak{gl}_{n}$-type monodromy matrix. In this section, we prove it by showing that the recurrence relations (4.21), coming from the $\mathfrak{o}_{2n+1}$ set-up, provide the recurrence relations (3.10), obtained in the $\mathfrak{gl}_{n}$ case. Obviously, to stay within the $\mathfrak{gl}_{n}$ part, we need to keep ${\bar{t}}^{0}=\varnothing$. Thus, we only consider the recurrence relations (4.21) with either $-n\leq\ell<k<0$ or $0<\ell<k\leq n$.
The case $0<\ell<k\leq n$ and ${\bar{t}}^{0}=\varnothing$. According to the first line in (4.16) it signifies that $|{\bar{t}}^{0}_{\scriptscriptstyle\rm I}|=0=\Theta(-i)+\Theta(-i-1)$ which is possible only for $i>0$. Since for these values of the index $i$ the step function $\Theta(-i-s-1)$ vanishes for any $s$, the cardinalities of the partitions (4.16) and (4.17) become the cardinalities (3.12). Moreover, for $i>0$, the sign $\delta_{-\ell-k}$ compensate the sign factor $\sigma_{-i}=-1$ and $[i]=i$ and $\ell_{i}=\ell$, so that the function $\Xi^{\ell,k}_{i,j}(z;{\bar{t}})$ in (4.20) becomes the function (3.11). The recurrence relation (4.21) takes the form
$\begin{split}&{\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}\}_{1}^{\ell-1},\{{\bar{t}}^{s},z\}_{\ell}^{k-1},\{{\bar{t}}^{s}\}_{k}^{n-1})=\\ &\quad=\sum_{i=1}^{\ell}\sum_{j=k}^{n}\sum_{\rm part}\frac{{\mathbb{T}}_{i,j}(z)\cdot{\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell-1},\{{\bar{t}}^{s}\}_{\ell}^{k-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{k}^{j-1},\{{\bar{t}}^{s}\}_{j}^{n-1})}{\lambda_{k}(z)\,g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II})\,h({\bar{t}}^{\ell},z)\,h(z,{\bar{t}}^{k-1})\,g({\bar{t}}^{k}_{\scriptscriptstyle\rm II},z)}\\ &\qquad\times\prod_{s=i}^{\ell-1}\Omega^{R}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})\prod_{s=k}^{j-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm I})\ \Omega^{L}({\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})\,,\end{split}$ (4.35)
where the cardinalities of the subsets in the sum over partitions are given by (3.12) and ${\mathbb{T}}_{i,j}(z)\equiv T_{i,j}(z)$ for $1\leq i,j\leq n$ as it was described for the first type of embedding of $Y(\mathfrak{gl}_{n})$ in $Y(\mathfrak{o}_{2n+1})$ (2.13). So we recover the $\mathfrak{gl}_{n}$-type recurrence relations (3.10).
The case $-n\leq\ell<k<0$ and ${\bar{t}}^{0}=\varnothing$. This case is more subtle because it corresponds to the off-shell Bethe vector $\hat{\mathbb{B}}({\bar{t}})$, symmetric to ${\mathbb{B}}({\bar{t}})$ in the sence of the paper [21]. If according to the recurrence relations (3.10) the off-shell Bethe vector ${\mathbb{B}}({\bar{t}})$ is built from the entries $T_{i,j}(z)$, $1\leq i<j\leq n$ of $\mathfrak{gl}_{n}$-type, the symmetric off-shell Bethe vector $\hat{\mathbb{B}}({\bar{t}})$ is built from the $\mathfrak{gl}_{n}$-type matrix entries ${\mathbb{T}}_{i,j}(z)$ described in the lemma 2.1. Let us consider the recurrence relations (4.21) at $-n\leq\ell<k<0$ and empty set ${\bar{t}}^{0}=\varnothing$ in more details.
If ${\bar{t}}^{0}=\varnothing$ then according to the first line in (4.17) $|{\bar{t}}^{0}_{\scriptscriptstyle\rm III}|=0=\Theta(j)+\Theta(j-1)$ which is possible only for strictly negative values of the index $j$. So the recurrence relation (4.21) becomes
$\begin{split}\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})\Big{|}_{{\bar{t}}^{0}=\varnothing}&={\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}\}_{1}^{-k-1},\{{\bar{t}}^{s},z_{s}\}_{-k}^{-\ell-1},\{{\bar{t}}^{s}\}_{-\ell}^{n-1})=\\ &=\mu^{k}_{\ell}(z;{\bar{t}})^{-1}\ \sum_{i=-n}^{\ell}\sum_{j=k}^{-1}\sum_{\rm part}\Xi^{\ell,k}_{i,j}(z;{\bar{t}})\ T_{i,j}(z)\cdot{\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{1}^{n-1})\,,\end{split}$ (4.36)
where $\mu^{k}_{\ell}(z;{\bar{t}})=\lambda_{k}(z)\,\psi_{\ell}(z;{\bar{t}})\,\phi_{k}(z;{\bar{t}})$ and
$\Xi^{\ell,k}_{i,j}(z;{\bar{t}})=\psi_{\ell}(z;{\bar{t}}_{\scriptscriptstyle\rm I})\ \phi_{k}(z;{\bar{t}}_{\scriptscriptstyle\rm III})\ \prod_{s=-\ell}^{-i-1}\Omega^{R}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{{\scriptscriptstyle\rm II}}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})\ \prod_{s=-j}^{-k-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm III})\Omega^{L}({\bar{t}}_{\scriptscriptstyle\rm II},{\bar{t}}_{{\scriptscriptstyle\rm III}}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm III},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})$ (4.37)
since ${\bar{t}}^{s}_{\scriptscriptstyle\rm III}=\varnothing$ for $-\ell\leq s\leq-i-1$. In (4.37) we dropped the index $s$ of the functions $\Omega^{L}_{s}$ and $\Omega^{R}_{s}$ because for positive $s$ these functions does not depend on $s$ and coincide with the functions (3.8).
The functions $\psi_{\ell}(z;{\bar{t}})$ and $\phi_{k}(z;{\bar{t}})$ are
$\psi_{\ell}(z;{\bar{t}})=\frac{g({\bar{t}}^{-\ell},z_{-l})}{g(z_{-l},{\bar{t}}^{-\ell-1})},\qquad\phi_{k}(z;{\bar{t}})=\frac{g(z_{-k-1},{\bar{t}}^{-k-1})}{g({\bar{t}}^{-k},z_{-k-1})}$ (4.38)
and the cardinalities of the partitions in (4.36) are
$|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=\begin{cases}0,&s<-\ell,\\ \Theta(-i-s-1),&s\geq-\ell,\end{cases}\qquad|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=\begin{cases}\Theta(j+s),&s<-k,\\ 0,&s\geq-k.\end{cases}$ (4.39)
Note again that there are no summations in the r.h.s. of (4.36) in case when $\ell=-n$ and $k=-1$.
Let us perform the following transformation of the recurrence relation (4.36). First we replace the strictly negative indices $-n\leq\ell<k<0$ by the strictly positive indices $0<{\ell^{\prime}}<{k^{\prime}}\leq n$ defined as
${\ell^{\prime}}=n+1+\ell,\quad 1\leq{\ell^{\prime}}\leq n-2,\qquad{k^{\prime}}=n+1+k,\quad 2\leq{k^{\prime}}\leq n.$ (4.40)
Analogously, we replace the summation over strictly negative indices $-n\leq i\leq\ell$ and $k\leq j\leq-1$ in (4.36) by the summation over strictly positive indices ${i^{\prime}}$ and ${j^{\prime}}$ such that
${i^{\prime}}=i+n+1,\quad 1\leq{i^{\prime}}\leq{\ell^{\prime}},\qquad{j^{\prime}}=j+n+1,\quad{k^{\prime}}\leq{j^{\prime}}\leq n\,.$ (4.41)
Moreover, instead of the sets of Bethe parameters ${\bar{t}}^{s}$ we consider the shifted sets of Bethe parameters
${\bar{\tau}}^{s}={\bar{t}}^{\,n-s}-c\,s=\{{\bar{t}}^{\,n-s}_{1}-c\,s,\ldots,{\bar{t}}^{\,n-s}_{r_{n-s}}-c\,s\},\qquad s=1,\ldots,n-1\,.$ (4.42)
These sets will be parted in the sum of the partitions ${\bar{\tau}}^{s}\vdash\{{\bar{\tau}}^{s}_{\scriptscriptstyle\rm I},{\bar{\tau}}^{s}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s}_{\scriptscriptstyle\rm III}\}$ with cardinalities according to (4.39) equal to
$|{\bar{\tau}}^{s}_{\scriptscriptstyle\rm I}|=\begin{cases}\Theta(s-{i^{\prime}}),&s<{\ell^{\prime}},\\ 0,&s\geq{\ell^{\prime}},\end{cases}\qquad|{\bar{\tau}}^{s}_{\scriptscriptstyle\rm III}|=\begin{cases}0,&s<{k^{\prime}},\\ \Theta({j^{\prime}}-s-1),&s\geq{k^{\prime}}.\end{cases}$ (4.43)
For $2\leq{k^{\prime}}\leq n$ we introduce the functions $\hat{\lambda}_{k^{\prime}}(u)$ and $\hat{\alpha}_{{k^{\prime}}}(u)$ by the equations
$\begin{split}\hat{\lambda}_{k^{\prime}}(u)&=\frac{1}{\lambda_{n+1-{k^{\prime}}}(u+c({k^{\prime}}-1))}\prod_{s=1}^{{k^{\prime}}-1}\frac{\lambda_{n+1-s}(u+cs)}{\lambda_{n+1-s}(u+c(s-1))},\\ \hat{\alpha}_{{k^{\prime}}}(u)&=\frac{\hat{\lambda}_{k^{\prime}}(u)}{\hat{\lambda}_{{k^{\prime}}+1}(u)}=\frac{\lambda_{n-{k^{\prime}}}(u+c{k^{\prime}})}{\lambda_{n+1-{k^{\prime}}}(u+c{k^{\prime}})}=\frac{1}{\alpha_{n-{k^{\prime}}}(u+c{k^{\prime}})}\,.\end{split}$ (4.44)
Then according to the relation (2.18) which connect the values of the functions $\lambda_{k}(u)$ for negative and positive values of the index $k$ and definition (4.42) one gets
$\lambda_{k}(z)=\hat{\lambda}_{k^{\prime}}(z_{n})\quad\mbox{and}\quad\hat{\alpha}_{s}({\bar{\tau}}^{s})=\alpha_{n-s}({\bar{t}}^{n-s})\,.$ (4.45)
Using all these definitions one can calculate
$\frac{1}{\psi_{\ell}(z;{\bar{t}}_{\scriptscriptstyle\rm II})}\prod_{s=-\ell}^{-i-1}\Omega^{R}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{{\scriptscriptstyle\rm II}}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})=\frac{(-1)^{|{\bar{\tau}}^{{i^{\prime}}-1}|-|{\bar{\tau}}^{{i^{\prime}}}_{\scriptscriptstyle\rm II}|}}{g(z_{n},{\bar{\tau}}^{{\ell^{\prime}}-1}_{\scriptscriptstyle\rm II})\,h({\bar{\tau}}^{{\ell^{\prime}}},z_{n})}\prod_{s={i^{\prime}}}^{{\ell^{\prime}}-1}\Omega^{R}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm I},{\bar{\tau}}^{s}_{{\scriptscriptstyle\rm II}}|{\bar{\tau}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s+1}_{\scriptscriptstyle\rm II})\,,$ (4.46)
$\begin{split}&\frac{1}{\phi_{k}(z;{\bar{t}}_{\scriptscriptstyle\rm II})}\prod_{s=-j}^{-k-1}\alpha_{s}({\bar{t}}^{s}_{\scriptscriptstyle\rm III})\Omega^{L}({\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{{\scriptscriptstyle\rm II}}|{\bar{t}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{s+1}_{\scriptscriptstyle\rm II})=\\ &\qquad=\frac{(-1)^{|{\bar{\tau}}^{{j^{\prime}}-1}_{\scriptscriptstyle\rm II}|-|{\bar{\tau}}^{{j^{\prime}}}|}}{h(z_{n},{\bar{\tau}}^{{k^{\prime}}-1})\,g({\bar{\tau}}^{{k^{\prime}}}_{\scriptscriptstyle\rm II},z_{n})}\prod_{s={k^{\prime}}}^{{j^{\prime}}-1}\hat{\alpha}_{s}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm III})\Omega^{L}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s}_{{\scriptscriptstyle\rm III}}|{\bar{\tau}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s+1}_{\scriptscriptstyle\rm II})\,.\end{split}$ (4.47)
and the recurrence relations (4.36) for the indices $\ell$ and $k$ such that $-n\leq\ell<k<0$ becomes
$\begin{split}&{\mathbb{B}}(\varnothing,\{{\bar{\tau}}^{s}\}_{1}^{{\ell^{\prime}}-1},\{{\bar{\tau}}^{s},z_{n}\}_{\ell^{\prime}}^{{k^{\prime}}-1},\{{\bar{\tau}}^{s}\}_{k^{\prime}}^{n-1})=\\ &=\sum_{{i^{\prime}}=1}^{\ell^{\prime}}\sum_{{j^{\prime}}={k^{\prime}}}^{n}\sum_{\rm part}\frac{T_{i,j}(z){(-1)^{{\sf h}_{i}+{\sf h}_{j}}}\cdot{\mathbb{B}}(\varnothing,\{{\bar{\tau}}^{s}\}_{1}^{{i^{\prime}}-1},\{{\bar{\tau}}^{s}_{\scriptscriptstyle\rm II}\}_{{i^{\prime}}}^{{\ell^{\prime}}-1},\{{\bar{\tau}}^{s}\}_{{\ell^{\prime}}}^{{k^{\prime}}-1},\{{\bar{\tau}}^{s}_{\scriptscriptstyle\rm II}\}_{{k^{\prime}}}^{{j^{\prime}}-1},\{{\bar{\tau}}^{s}\}_{{j^{\prime}}}^{n-1})}{\hat{\lambda}_{k^{\prime}}(z_{n})\,g(z_{n},{\bar{\tau}}^{{\ell^{\prime}}-1}_{\scriptscriptstyle\rm II})\,h({\bar{\tau}}^{\ell^{\prime}},z_{n})\,h(z_{n},{\bar{\tau}}^{{k^{\prime}}-1})\,g({\bar{\tau}}^{k^{\prime}}_{\scriptscriptstyle\rm II},z_{n})}\\ &\qquad\times\prod_{s={i^{\prime}}}^{{\ell^{\prime}}-1}\Omega^{R}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm I},{\bar{\tau}}^{s}_{\scriptscriptstyle\rm II}|{\bar{\tau}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s+1}_{\scriptscriptstyle\rm II})\prod_{s={k^{\prime}}}^{{j^{\prime}}-1}\hat{\alpha}_{s}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm III})\ \Omega^{L}({\bar{\tau}}^{s}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s}_{\scriptscriptstyle\rm III}|{\bar{\tau}}^{s-1}_{\scriptscriptstyle\rm II},{\bar{\tau}}^{s+1}_{\scriptscriptstyle\rm II})\,.\end{split}$ (4.48)
Due to the definition (2.21) and the relations (2.18) and (2.9), the generators ${\sf h}_{i}$ for $i<0$ are equal to $-{\sf h}_{-i}$ and according to the formulas (2.28) the action of the generators ${\sf h}_{i}$ for $-n\leq i\leq-1$ to the off-shell Bethe vectors is given by
${{\sf h}_{i}}\cdot{\mathbb{B}}({\bar{t}})=(|{\bar{t}}^{-i}|-|{\bar{t}}^{-i-1}|)\ {\mathbb{B}}({\bar{t}})=(|{\bar{t}}^{\,n+1-{i^{\prime}}}|-|{\bar{t}}^{\,n-{i^{\prime}}}|)\ {\mathbb{B}}({\bar{t}})=(|{\bar{\tau}}^{{i^{\prime}}-1}|-|{\bar{\tau}}^{{i^{\prime}}}|)\ {\mathbb{B}}({\bar{\tau}})\,.$ (4.49)
This explains the appearance of the operator $(-1)^{{\sf h}_{i}+{\sf h}_{j}}$ in the r.h.s. of (4.48). Due to (4.49) this operator generate the sign factors $(-1)^{|{\bar{\tau}}^{{i^{\prime}}-1}|-|{\bar{\tau}}^{{i^{\prime}}}_{\scriptscriptstyle\rm II}|}$ and $(-1)^{|{\bar{\tau}}^{{j^{\prime}}-1}_{\scriptscriptstyle\rm II}|-|{\bar{\tau}}^{{j^{\prime}}}|}$ which comes from (4.46) and (4.47) respectively.
Let us define the monodromy matrix entries ${\mathbb{T}}_{{i^{\prime}},{j^{\prime}}}(z)$ for all ${i^{\prime}},{j^{\prime}}=1,\ldots,n$ as:
${\mathbb{T}}_{{i^{\prime}},{j^{\prime}}}(z_{n})=-T_{i,j}(z)\ (-1)^{\delta_{i,j}+{{\sf h}_{i}+{\sf h}_{j}}},\qquad\text{with }\begin{cases}i={i^{\prime}}-n-1\,,\\ j={j^{\prime}}-n-1\,.\end{cases}$ (4.50)
Lemma 2.1 ensures that this monodromy matrix is of $\mathfrak{gl}_{n}$-type and satisfies the $RTT$ relation (2.4) with the $\mathfrak{gl}_{n}$-type $R$-matrix. The off-shell Bethe vectors $\hat{\mathbb{B}}({\bar{t}})$ which can be constructed by the $\mathfrak{gl}_{n}$-type monodromy (4.50) and the recurrence relations (4.48) are related to the original vectors ${\mathbb{B}}({\bar{t}})$ as follows [21]
$\hat{\mathbb{B}}(\varnothing,\{{\bar{t}}^{s}\}_{1}^{n-1})=(-1)^{|{\bar{\tau}}|}\ {\mathbb{B}}(\varnothing,\{{\bar{\tau}}^{s}\}_{1}^{n-1})=(-1)^{|{\bar{t}}|}\ {\mathbb{B}}(\varnothing,\{{\bar{t}}^{n-s}-c\,s\}_{1}^{n-1}),$ (4.51)
where $|{\bar{t}}|=|{\bar{\tau}}|$ is a total cardinality of the sets of Bethe parameters
$|{\bar{t}}|=\sum_{s=0}^{n-1}|{\bar{t}}^{s}|\,.$ (4.52)
The sign factor $(-1)^{|{\bar{t}}|}$ appears in the relation (4.51) because of the sign we introduced in definition (4.50) of the monodromy matrix ${\mathbb{T}}(z)$.
One can see now that the recurrence relation (4.36) or (4.48) for the Bethe vectors ${\mathbb{B}}({\bar{t}})$ coincides exactly with the recurrence relations (4.35) for the Bethe vectors $\hat{\mathbb{B}}({\bar{t}})$ if we remove the primes of the positive integers ${\ell^{\prime}},{k^{\prime}},{i^{\prime}},{j^{\prime}}$, replace $z_{n}\to z$, ${\bar{\tau}}^{s}\to{\bar{t}}^{s}$, $\hat{\lambda}_{k}(z_{n})\to\lambda_{k}(z)$, ${\mathbb{T}}_{i,j}(z_{n})\to{\mathbb{T}}_{i,j}(z)$. This proves that the recurrence relation (4.36) for the original Bethe vectors ${\mathbb{B}}({\bar{t}})$ is equivalent to the recurrence relation (4.35) for the Bethe vector $\hat{\mathbb{B}}({\bar{t}})$ (4.51) built from the monodromy matrix entries ${\mathbb{T}}_{i,j}(z)$ (4.50).
4.4 Embedding $Y(\mathfrak{o}_{2a+1})\otimes Y(\mathfrak{gl}_{n-a})\hookrightarrow Y(\mathfrak{o}_{2n+1})$
In this section we study the recurrence relations (4.21) in the case when the set ${\bar{t}}^{a}=\varnothing$ for some $a>0$.
Proposition 4.1.
The off-shell Bethe vectors $\mathbb{B}_{\mathfrak{o}}(\{{\bar{t}}^{s}\}_{0}^{a-1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})$ factorizes into the product of $\mathfrak{o}_{2a+1}$-type pre-Bethe vector $\mathcal{B}_{\mathfrak{o}}(\{{\bar{t}}^{s}\}_{0}^{a-1})$ and $\mathfrak{gl}_{n-a}$-type pre-Bethe vectors $\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})$:
$\mathbb{B}_{\mathfrak{o}}(\{{\bar{t}}^{s}\}_{0}^{a-1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})=\mathcal{B}_{\mathfrak{o}}(\{{\bar{t}}\}_{0}^{a-1})\cdot\mathcal{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})\,|0\rangle\,.$ (4.53)
Proof.
The proof of this proposition is analogous to the proof of the proposition 3.2. We consider the recurrence relations (4.21) for three different ranges of the indices $\ell$ and $k$: $-a\leq\ell<k\leq a$, $a+1\leq\ell<k\leq n$ and $-n\leq\ell<k\leq-a-1$. These ranges for $k$ and $\ell$ ensure that the set ${\bar{t}}^{a}$ remains empty. Since $\mathbb{B}_{\mathfrak{o}}(\varnothing)=\mathbb{B}_{\mathfrak{gl}}(\varnothing)=|0\rangle$, the recurrence relations will prove the statement of the proposition by induction on the cardinalities.
  • $-a\leq\ell<k\leq a$. According to (4.16), the cardinality of the empty subset ${\bar{t}}^{a}_{\scriptscriptstyle\rm I}$ is given by the step function $\Theta(-i-a-1)$, which must vanish. It immediately results that the summation in the recurrence relation (4.21) is restricted to the range $-a\leq i\leq\ell$. Analogously, the cardinality of the empty subset ${\bar{t}}^{a}_{\scriptscriptstyle\rm III}$ given by (4.17) is equal to $\Theta(j-a-1)$ which restricts the range in the summation over $j$ to $k\leq j\leq a$. This means that the recurrence relation for $\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})|_{{\bar{t}}^{a}=\varnothing}$ implies only the monodromy entries $T_{i,j}(z)$ for $-a\leq i\leq\ell$ and $k\leq j\leq a$. The recurrence relation $\mathcal{Z}^{a}_{-a}\cdot{\mathbb{B}}({\bar{t}})|_{{\bar{t}}^{a}=\varnothing}$ will have only one term corresponding to the action of monodromy entry $T_{-a,a}(z)$ on the off-shell Bethe vector.
    Moreover, the action of the zero mode operators ${\sf T}_{\ell+1,\ell}$ for $\ell=0,1,\ldots,a-1$ onto the Bethe vector (4.53)
    $\begin{split}{\sf T}_{\ell+1,\ell}\cdot\mathbb{B}_{\mathfrak{o}}&(\{{\bar{t}}^{s}\}_{0}^{a-1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})=\sum_{\rm part}\Big{(}\chi_{\ell+1}\ \alpha_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\,\Omega^{L}_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})\\ &-\chi_{\ell}\ \Omega^{R}_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})\Big{)}\mathbb{B}_{\mathfrak{o}}(\{{\bar{t}}^{s}\}_{0}^{\ell-1},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},\{{\bar{t}}^{s}\}_{\ell+1}^{a-1},\varnothing,\{{\bar{t}}^{s}\}_{a+1}^{n-1})\end{split}$ (4.54)
    does not affect the sets of Bethe parameters $\{{\bar{t}}^{s}\}_{a+1}^{n-1}$ and so coincides with the action of the zero mode operators onto the Bethe vectors in $\mathfrak{o}_{2a+1}$-invariant model. This shows that the recurrence relations for $\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})|_{{\bar{t}}^{a}=\varnothing}$ for $-a\leq\ell<k\leq a$ coincides with the $\mathfrak{o}_{2a+1}$ recurrence relation for the vector $\mathbb{B}_{\mathfrak{o}}(\{{\bar{t}}\}_{0}^{a-1})$.
  • $a+1\leq\ell<k\leq n$. In this case the cardinalities of the empty subsets ${\bar{t}}^{a}_{\scriptscriptstyle\rm I}={\bar{t}}^{a}_{\scriptscriptstyle\rm III}=\varnothing$ given by the equalities (4.16) and (4.17) yield a restriction only for the index $i\geq a+1$. The summation over $i$ and $j$ in the r.h.s. of the recurrence relation for $\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})|_{{\bar{t}}^{a}=\varnothing}$ in this case will be restricted to the range $a+1\leq i\leq\ell<k\leq j\leq n$. The Bethe vector $\mathcal{Z}^{n}_{a+1}\cdot{\mathbb{B}}({\bar{t}})$ will be proportional to the action of the monodromy matrix entry $T_{a+1,n}(z)\cdot{\mathbb{B}}({\bar{t}})$ and the action of the zero mode operators ${\sf T}_{\ell+1,\ell}$ for $\ell=a+1,\ldots,n-1$ onto Bethe vectors (4.53) coincide with $\mathfrak{gl}_{n-a}$-type actions (3.7). The recurrence relation (4.21) in this case will not affect the sets of Bethe parameters $\{{\bar{t}}^{s}\}_{0}^{a-1}$ and will coincides with the $\mathfrak{gl}_{n-a}$-type recurrence relation for the Bethe vector $\mathcal{Z}^{k}_{\ell}\cdot\mathbb{B}_{\mathfrak{gl}}(\{{\bar{t}}^{s}\}_{a+1}^{n-1})$, see (4.35).
  • $-n\leq\ell<k\leq-a-1$. Using again (4.16) and (4.17) we can observe that the summation over $i$ and $j$ in the r.h.s. of the recurrence relation for the Bethe vector $\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})|_{{\bar{t}}^{a}=\varnothing}$ will be restricted to the range $-n\leq i\leq\ell<k\leq j\leq-a-1$. Using arguments of the section 4.3 one can show that the corresponding recurrence relation for $-n\leq\ell<k\leq-a-1$ will be equivalent to the $\mathfrak{gl}_{n-a}$-type recurrence relation for $\hat{\mathbb{B}}({\bar{t}})$, where $\hat{\mathbb{B}}({\bar{t}})$ is the Bethe vector symmetric to ${\mathbb{B}}({\bar{t}})$ in the sense of the symmetry introduced in the paper [21].
$\square$
Remark 4.2.
When $a=n-1$ the recurrence relations described in the second and third items are absent and the recurrence relations described in the first item yield the reduction over rank $n\to n-1$ of the recurrence relations for the off-shell Bethe vectors in $\mathfrak{o}_{2n-1}$-invariant integrable models.
Remark 4.3 (Splitting property).
Using the statement of the proposition 4.1 and the rectangular recurrence relations (4.21) for the off-shell Bethe vectors in $\mathfrak{o}_{2n+1}$-invariant integrable models one can prove a splitting property for these vectors, similar to (3.15). Again, the splitting property is a direct consequence of the presentation of the Bethe vectors within the projection method [18].
5 Conclusion
Using the zero modes method we describe in this paper the rectangular recurrence relations for the off-shell Bethe vectors in $\mathfrak{gl}_{n}$- and $\mathfrak{o}_{2n+1}$-invariant integrable models. These relations are presented as sums of the actions of the monodromy entries $T_{i,j}(z)$ with $i<j$ on the off-shell Bethe vectors ${\mathbb{B}}({\bar{t}})$ and sums over partitions of the sets of Bethe parameters. When one of the set ${\bar{t}}^{s}$ becomes empty the recurrence relations given by theorem 4.1 describe a factorization of the Bethe vectors and relates to the symmetry of $\mathfrak{gl}_{n}$-type Bethe vectors found in the paper [21]. Whether there exists a similar symmetry for $\mathfrak{o}_{2n+1}$-type Bethe vectors remains an open question. Such a symmetry would allow to relate the recurrence relation (4.24) to the recurrence relation (4.22). If it exists, this symmetry would require to explore the explicit relations between the monodromy matrix entries $T_{i,j}(z)$ following from the symmetry relation (2.9). It makes the search for a symmetry relation of $\mathfrak{o}_{2n+1}$-type monodromy matrix much more involved than in the $\mathfrak{gl}_{n}$ case.
Now that the recurrence relations are obtained, one can deduce recurrence relations for scalar products, and for their building block, the so-called highest coefficients. One can also compute the norm of the on-shell Bethe vectors and express it as a determinant. Calculations are under review, and the results will appear soon.
Finally, the outcome obtained in this paper can be generalized for the Bethe vectors in $\mathfrak{sp}_{2n}$- and $\mathfrak{so}_{2n}$-invariant integrable models. The procedure is similar to the one presented here and is currently under investigation. The issue will be published elsewhere.
Acknowledgement
We are grateful to Alexander Molev for fruitful discussions on embeddings in Yangian algebras. S.P. acknowledges the support of the PAUSE Programme and hospitality at LAPTh where this work was done. The research of A.L. was supported by Beijing Natural Science Foundation (IS24006). A.L. is also grateful to the CNRS PHYSIQUE for support during his visit to Annecy in the course of this investigation.
Appendix A Proof of the proposition 3.1
The proof of this proposition is based on the following lemma.
Lemma A.1.
We recall that $\bar{u}^{s}=\{\bar{t}^{s},z\}$. The off-shell Bethe vector ${\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},\{\bar{u}^{s}\}_{\ell}^{n-1})$ for $1\leq\ell<n$ can be presented in the form
$\begin{split}\mathcal{Z}_{\ell}^{n}\cdot{\mathbb{B}}({\bar{t}})&=\frac{1}{\mu_{\ell}^{n}(z;{\bar{t}})}\ \sum_{i=1}^{\ell}\sum_{\rm part}g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\ \prod_{p=i}^{\ell-1}\Omega^{R}({\bar{t}}^{p}_{\scriptscriptstyle\rm I},{\bar{t}}^{p}_{\scriptscriptstyle\rm II}|{\bar{t}}^{p-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{p+1}_{\scriptscriptstyle\rm II})\times\\ &\quad\times T_{i,n}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell-1},\{{\bar{t}}^{s}\}_{\ell}^{n-1})\,,\end{split}$ (A.1)
where the sum goes over partitions ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm I},{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}$ with cardinalities $|{\bar{t}}^{s}_{\scriptscriptstyle\rm I}|=1$ for all $s=i,\ldots,\ell-1$ while ${\bar{t}}^{\ell}$ is not parted with ${\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\equiv{\bar{t}}^{\ell}$.
Proof.
We prove relation (A.1) through an induction over $\ell$. Equality (3.5) can be considered as the base of this induction since for $\ell=1$ there are no partitions in the r.h.s. of (A.1) and it becomes identical to (3.5).
Assuming that equality (A.1) is valid for some fixed $\ell<n-1$, we multiply it by the normalization factor $\mu^{n}_{\ell}(z;{\bar{t}})$ and apply to both sides the zero mode operator ${\sf T}_{\ell+1,\ell}$. Using the action of the zero modes (3.7), we get from the l.h.s. of (A.1):
$\begin{split}&-\chi_{\ell}\ \mu^{n}_{\ell+1}(z;{\bar{t}})\ \mathcal{Z}_{\ell+1}^{n}\cdot{\mathbb{B}}({\bar{t}})\ +\\ &\qquad+\sum_{\rm part}\Big{(}\chi_{\ell+1}\,\alpha_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\,\Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})-\chi_{\ell}\ f({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)\ \Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})\Big{)}\times\\ &\qquad\times\mu^{n}_{\ell}(z;{\bar{t}}^{\ell-1},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{n-1})\ \mathcal{Z}_{\ell}^{n}\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},\{{\bar{t}}^{s}\}_{\ell+1}^{n-1})\,.\end{split}$ (A.2)
Using the induction assumption, relation (A.2) rewrites
$\begin{split}&-\chi_{\ell}\ \mu^{n}_{\ell+1}(z;{\bar{t}})\ \mathcal{Z}_{\ell+1}^{n}\cdot{\mathbb{B}}({\bar{t}})\ +\\ &\qquad+\sum_{\rm part}\Big{(}\chi_{\ell+1}\,\alpha_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\,\Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})-\chi_{\ell}\ f({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)\ \Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})\Big{)}\times\\ &\qquad\times\sum_{i=1}^{\ell}g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\ \prod_{p=i}^{\ell-1}\Omega^{R}({\bar{t}}^{p}_{\scriptscriptstyle\rm I},{\bar{t}}^{p}_{\scriptscriptstyle\rm II}|{\bar{t}}^{p-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{p+1}_{\scriptscriptstyle\rm II})\ T_{i,n}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell},\{{\bar{t}}^{s}\}_{\ell+1}^{n-1})\,.\end{split}$ (A.3)
To get (A.2) and (A.3) we have used the equalities
$\begin{split}\Omega^{L}({\bar{t}}^{\ell},z|{\bar{t}}^{\ell-1},\bar{u}^{\ell+1})&=0\,,\\ \Omega^{R}(z,{\bar{t}}^{\ell}|{\bar{t}}^{\ell-1},\bar{u}^{\ell+1})&=\frac{g(z,{\bar{t}}^{\ell})}{h({\bar{t}}^{\ell},z)}\ \frac{h({\bar{t}}^{\ell+1},z)}{g(z,{\bar{t}}^{\ell-1})}\,,\\ \Omega^{L}(\{{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},z\},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1},\bar{u}^{\ell+1})&=\frac{1}{h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)}\ \Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})\,,\\ \Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},\{z,{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\}|{\bar{t}}^{\ell-1},\bar{u}^{\ell+1})&=g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)\ \Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1},{\bar{t}}^{\ell+1})\,,\\ \mu^{n}_{\ell}(z;{\bar{t}})\ \frac{g(z,{\bar{t}}^{\ell})}{h({\bar{t}}^{\ell},z)}\ \frac{h({\bar{t}}^{\ell+1},z)}{g(z,{\bar{t}}^{\ell-1})}&=\mu^{n}_{\ell+1}(z;{\bar{t}})\,,\\ \mu^{n}_{\ell}(z;{\bar{t}})&=h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)\ \mu^{n}_{\ell}(z;{\bar{t}}^{\ell-1},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{n-1})\,.\end{split}$
Using the commutation relations
$[{\sf T}_{\ell+1,\ell},T_{i,n}(z)]=-\chi_{\ell}\ \delta_{\ell,i}\ T_{\ell+1,n}(z)$ (A.4)
the action of the zero mode ${\sf T}_{\ell+1,\ell}$ on the r.h.s. of the induction assumption (A.1) reads
$\begin{split}-\chi_{\ell}\ T_{\ell+1,n}(z)\cdot{\mathbb{B}}({\bar{t}})&+\sum_{i=1}^{\ell}\sum_{\rm part}g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\prod_{p=i}^{\ell-1}\Omega^{R}({\bar{t}}^{p}_{\scriptscriptstyle\rm I},{\bar{t}}^{p}_{\scriptscriptstyle\rm II}|{\bar{t}}^{p-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{p+1}_{\scriptscriptstyle\rm II})\Big{|}_{{\bar{t}}^{\ell}\vdash\{{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\}}\times\\ &\times\Big{(}\chi_{\ell+1}\,\alpha_{\ell}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\,\Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell+1})-\chi_{\ell}\ \Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell+1})\Big{)}\times\\ &\qquad\times T_{i,n}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell},\{{\bar{t}}^{s}\}_{\ell+1}^{n-1})\Big{)}\,.\end{split}$ (A.5)
Let us compare the coefficients of the twisting parameters $\chi_{\ell+1}$ and $\chi_{\ell}$ in (A.3) and (A.5). Due to the relations
$\begin{split}&\Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|\{{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II}\},{\bar{t}}^{\ell+1})=h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\ \Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell+1})\\ &\Omega^{R}({\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-2}_{\scriptscriptstyle\rm II},\{{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}\})=h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\ \Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I}|{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell+1})\end{split}$
the terms proportional to $\chi_{\ell+1}$ in (A.3) and in (A.5) are equal, and cancel each other when we equate (A.3) and (A.5). Then $\chi_{\ell}$ factorizes globally, and we get a relation with no explicit dependence of the $\chi_{i}$ parameters: this is a manifestation of the principle described in remark 2.2.
On the other hand using in (A.3) the equality
$\Omega^{R}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|\{{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II}\},{\bar{t}}^{\ell+1})=\frac{1}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}\ \Omega^{L}({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell}_{\scriptscriptstyle\rm II}|{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{\ell+1})$
we conclude that the equality between (A.3) and (A.5) is equivalent to the relation (A.1) for $\ell\to\ell+1$ due to the identity
$\begin{split}&g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\left(h({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})-\frac{f({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}\right)=\frac{g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}\Big{(}f({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})-f({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)\Big{)}=\\ &\quad=\frac{g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})}\Big{(}g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})-g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)\Big{)}=\frac{g(z,{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\,g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\,g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},z)}{g({\bar{t}}^{\ell}_{\scriptscriptstyle\rm I},{\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I})\,g({\bar{t}}^{\ell-1}_{\scriptscriptstyle\rm I},z)}=g(z,{\bar{t}}^{\ell}_{\scriptscriptstyle\rm I})\,.\end{split}$ (A.6)
This ends the inductive proof of relation (A.1). $\square$
End of proposition’s proof.
To finish the proof of the proposition 3.1 one has to perform an inductive proof over $k$ for the recurrence relation (3.10). We will consider the induction step $k+1\to k$ taking as induction base the just proved equality (A.1).
Let us assume that equality (3.10) is valid for $\mathcal{Z}_{\ell}^{k+1}$ for some $k<n$. The induction proof means that this assumption should lead to the equality (3.10) for $\mathcal{Z}_{\ell}^{k}$. To perform the induction step we multiply both sides of (3.10) at $k+1$ by the function $\mu^{k+1}_{\ell}(z;{\bar{t}})$ and act on this relation by the zero mode operator ${\sf T}_{k+1,k}$. Using the equalities
$\begin{split}\Omega^{L}({\bar{t}}^{k},z|\bar{u}^{k-1},{\bar{t}}^{k+1})&=\frac{g({\bar{t}}^{k},z)}{h(z,{\bar{t}}^{k})}\ \frac{h(z,{\bar{t}}^{k-1})}{g({\bar{t}}^{k+1},z)}\,,\\ \Omega^{R}(z,{\bar{t}}^{k}|\bar{u}^{k-1},{\bar{t}}^{k+1})&=0\,,\\ \Omega^{L}(\{{\bar{t}}^{k}_{\scriptscriptstyle\rm II},z\},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}|\bar{u}^{k-1},{\bar{t}}^{k+1})&=g(z,{\bar{t}}^{k}_{\scriptscriptstyle\rm III})\ \Omega^{L}({\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}|{\bar{t}}^{k-1},{\bar{t}}^{k+1})\,,\\ \Omega^{R}({\bar{t}}^{k}_{\scriptscriptstyle\rm III},\{{\bar{t}}^{k}_{\scriptscriptstyle\rm II},z\}|\bar{u}^{k-1},{\bar{t}}^{k+1})&=\frac{1}{h(z,{\bar{t}}^{k}_{\scriptscriptstyle\rm III})}\ \Omega^{R}({\bar{t}}^{k}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm II}|{\bar{t}}^{k-1},{\bar{t}}^{k+1})\,,\\ \alpha_{k}(z)\ \mu^{k+1}_{\ell}(z;{\bar{t}})\ \frac{g({\bar{t}}^{k},z)}{h(z,{\bar{t}}^{k})}\ \frac{h(z,{\bar{t}}^{k-1})}{g({\bar{t}}^{k+1},z)}&=\mu^{k}_{\ell}(z;{\bar{t}})\,,\\ \mu^{k+1}_{\ell}(z;{\bar{t}})&=h(z,{\bar{t}}^{k}_{\scriptscriptstyle\rm III})\ \mu^{k+1}_{\ell}(z;{\bar{t}}^{l-1},{\bar{t}}^{\ell},{\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k+1})\end{split}$
one can write the l.h.s. of the resulting relation as as follows
$\begin{split}&\mu^{k+1}_{\ell}(z;{\bar{t}})\ {\sf T}_{k+1,k}\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{\ell-1},\{\bar{u}^{s}\}_{\ell}^{k},\{{\bar{t}}^{s}\}_{k+1}^{n-1})=\chi_{k+1}\,\mu^{k}_{\ell}(z;{\bar{t}})\,\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})\ +\\ &\quad+\sum_{i=1}^{\ell}\sum_{j=k+1}^{n}\sum_{\rm part}\Big{(}\chi_{k+1}\,\alpha_{k}({\bar{t}}^{k}_{\scriptscriptstyle\rm III})\,\Omega^{L}({\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}|{\bar{t}}^{k-1},{\bar{t}}^{k+1})\,f(z,{\bar{t}}^{k}_{\scriptscriptstyle\rm III})-\chi_{k}\,\Omega^{R}({\bar{t}}^{k}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm II}|{\bar{t}}^{k-1},{\bar{t}}^{k+1})\Big{)}\times\\ &\quad\times\Xi^{\ell,k+1}_{i,j}(z;{\bar{t}})\ T_{i,j}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell-1},\{{\bar{t}}^{s}\}_{\ell}^{k-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{k}^{j-1},\{{\bar{t}}^{s}\}_{j}^{n-1})\,.\end{split}$ (A.7)
Here for further convenience, the partition resulting from the action of the zero mode operator ${\sf T}_{k+1,k}$ is noted ${\bar{t}}^{k}\vdash\{{\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}\}$ with cardinality $|{\bar{t}}^{k}_{\scriptscriptstyle\rm III}|=1$.
On the other hand the r.h.s. of the same equality takes the form
$\begin{split}&\chi_{k+1}\,\sum_{i=1}^{\ell}\sum_{\rm part}\Xi^{\ell,k+1}_{i,k+1}(z,{\bar{t}})\ T_{i,k}(z)\cdot{\mathbb{B}}({\bar{t}}_{\scriptscriptstyle\rm II})+\sum_{i=1}^{\ell}\sum_{j=k+1}^{n}\sum_{\rm part}\Xi^{\ell,k+1}_{i,j}(z;{\bar{t}})\Big{|}_{{\bar{t}}^{k}\vdash\{{\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}\}}\times\\ &\qquad\times\Big{(}\chi_{k+1}\,\alpha_{k}({\bar{t}}^{k}_{\scriptscriptstyle\rm III})\,\Omega^{L}({\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}|{\bar{t}}^{k-1},{\bar{t}}^{k+1}_{\scriptscriptstyle\rm II})\,-\chi_{k}\,\Omega^{R}({\bar{t}}^{k}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm II}|{\bar{t}}^{k-1},{\bar{t}}^{k+1}_{\scriptscriptstyle\rm II})\Big{)}\times\\ &\qquad\times T_{i,j}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{i-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{i}^{\ell-1},\{{\bar{t}}^{s}\}_{\ell}^{k-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{k}^{j-1},\{{\bar{t}}^{s}\}_{j}^{n-1})\,,\end{split}$ (A.8)
where we used the commutation relation (3.6) and took into account that the summation index $i$ in (A.8) satisfies the inequalities $1\leq i\leq\ell<k$. Due to the formulas
$\begin{split}\Xi^{\ell,k+1}_{i,k+1}(z;{\bar{t}})=\Xi^{\ell,k}_{i,k}(z;{\bar{t}})&=\psi_{\ell}(z;{\bar{t}}_{\scriptscriptstyle\rm I})\prod_{p=i}^{\ell-1}\Omega^{R}({\bar{t}}^{p}_{\scriptscriptstyle\rm I},{\bar{t}}^{p}_{\scriptscriptstyle\rm II}|{\bar{t}}^{p-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{p+1}_{\scriptscriptstyle\rm II})\,,\\ \Xi^{\ell,k+1}_{i,j}(z;{\bar{t}})\Big{|}_{{\bar{t}}^{k}\vdash\{{\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}\}}&=h({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})\ \Xi^{\ell,k+1}_{i,j}(z;{\bar{t}})\Big{|}_{{\bar{t}}^{k}\vdash\{{\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}\}}\,,\\ \Omega^{R}({\bar{t}}^{k}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm II}|{\bar{t}}^{k-1},{\bar{t}}^{k+1})\Big{|}_{{\bar{t}}^{k+1}\vdash\{{\bar{t}}^{k+1}_{\scriptscriptstyle\rm II},{\bar{t}}^{k+1}_{\scriptscriptstyle\rm III}\}}&=h({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})\ \Omega^{R}({\bar{t}}^{k}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm II}|{\bar{t}}^{k-1},{\bar{t}}^{k+1}_{\scriptscriptstyle\rm II})\,,\\ \Omega^{L}({\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}|{\bar{t}}^{k-1},{\bar{t}}^{k+1})\Big{|}_{{\bar{t}}^{k+1}\vdash\{{\bar{t}}^{k+1}_{\scriptscriptstyle\rm II},{\bar{t}}^{k+1}_{\scriptscriptstyle\rm III}\}}&=\frac{1}{g({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})}\ \Omega^{L}({\bar{t}}^{k}_{\scriptscriptstyle\rm II},{\bar{t}}^{k}_{\scriptscriptstyle\rm III}|{\bar{t}}^{k-1},{\bar{t}}^{k+1}_{\scriptscriptstyle\rm II})\end{split}$
and the identity for the rational functions
$\phi_{k+1}(z;{\bar{t}}_{\scriptscriptstyle\rm III})\Big{(}h({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})-\frac{f(z,{\bar{t}}^{k}_{\scriptscriptstyle\rm iii})}{g({\bar{t}}^{k+1}_{\scriptscriptstyle\rm III},{\bar{t}}^{k}_{\scriptscriptstyle\rm III})}\Big{)}=\phi_{k}(z,{\bar{t}}_{\scriptscriptstyle\rm III})$
the equality between (A.7) and (A.8) is equivalent to (3.10). This finishes the inductive proof of the proposition 3.1. $\square$
Alternative proof.
Proposition 3.1 can also be proven using an induction $\ell\to\ell+1$ and taking as induction base the relation (A.9), coming from following alternative lemma.
Lemma A.2.
The off-shell Bethe vector $\mathcal{Z}_{1}^{k}\cdot{\mathbb{B}}({\bar{t}})={\mathbb{B}}(\{\bar{u}^{s}\}_{1}^{k-1},\{{\bar{t}}^{s}\}_{k}^{n-1})$ for $1<k\leq n$ can be presented in the form
$\begin{split}\mathcal{Z}_{1}^{k}\cdot{\mathbb{B}}({\bar{t}})&=\frac{1}{\mu_{1}^{k}(z;{\bar{t}})}\ \sum_{j=k}^{n}\sum_{\rm part}g({\bar{t}}^{k}_{\scriptscriptstyle\rm III},z)\ \prod_{p=k}^{j-1}\alpha_{p}({\bar{t}}^{p}_{\scriptscriptstyle\rm III})\ \Omega^{L}({\bar{t}}^{p}_{\scriptscriptstyle\rm II},{\bar{t}}^{p}_{\scriptscriptstyle\rm III}|{\bar{t}}^{p-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{p+1}_{\scriptscriptstyle\rm II})\times\\ &\quad\times T_{1,j}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{1}^{k-1},\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II}\}_{k}^{j-1},\{{\bar{t}}^{s}\}_{j}^{n-1})\end{split}$ (A.9)
where the sum goes over partitions ${\bar{t}}^{s}\vdash\{{\bar{t}}^{s}_{\scriptscriptstyle\rm II},{\bar{t}}^{s}_{\scriptscriptstyle\rm III}\}$ with cardinalities $|{\bar{t}}^{s}_{\scriptscriptstyle\rm III}|=1$ for all $s=k,\ldots,j-1$ while the set ${\bar{t}}^{k-1}$ is not parted and ${\bar{t}}^{k-1}_{\scriptscriptstyle\rm II}\equiv{\bar{t}}^{k-1}$.
The proofs of the lemma A.2 and the end of recursion for proposition 3.1 are similar as above, and we don’t reproduce them in the present paper.
Appendix B Sketch of the proof of the theorem 4.1
The proof of this theorem follows the method described in the appendix A. The starting point is the simple recurrence relation (4.6) for the Bethe vector ${\mathbb{B}}(\{\bar{w}^{s}\}_{0}^{n-1})$. Applying the zero mode operator ${\sf T}_{n,n-1}$ to this relation and using the commutation relations between the zero modes and the monodromy entries (4.7) as well as the action of the zero mode operators on the off-shell Bethe vectors (4.8) we get a relation which involves terms proportional to $\chi_{n}$ and terms proportional to $\chi_{n-1}$. Since the relation involves only Bethe vectors, monodromy matrix entries $T_{ij}(z)$ and eigenvalues $\lambda_{i}(z)$, due to remark 2.2, the coefficients of these two independent twisting parameters yield two recurrence relations, one for the Bethe vectors ${\mathbb{B}}(\{\bar{w}^{s}\}_{0}^{n-2},\bar{u}^{n-1})=\mathcal{Z}^{n}_{-n+1}\cdot{\mathbb{B}}({\bar{t}})$ and one for the Bethe vector ${\mathbb{B}}(\{\bar{w}^{s}\}_{0}^{n-2},\bar{v}^{n-1})=\mathcal{Z}^{n-1}_{-n}\cdot{\mathbb{B}}({\bar{t}})$. These relations are
$\begin{split}\mathcal{Z}_{-n+1}^{n}\cdot{\mathbb{B}}({\bar{t}})=\frac{1}{\mu_{-n+1}^{n}(z;{\bar{t}})}\sum_{i=-n}^{-n+1}\sum_{\rm part}&g({\bar{t}}^{n-1}_{\scriptscriptstyle\rm I},z_{n-1})\,\Omega^{R}({\bar{t}}^{n-1}_{\scriptscriptstyle\rm I},{\bar{t}}^{n-1}_{\scriptscriptstyle\rm II}|{\bar{t}}^{n-2},\varnothing)\times\\ &\quad\times T_{i,n}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{0}^{n-2},{\bar{t}}^{n-1}_{\scriptscriptstyle\rm II})\,,\end{split}$ (B.1)
$\begin{split}\mathcal{Z}_{-n}^{n-1}\cdot{\mathbb{B}}({\bar{t}})=\frac{1}{\mu_{-n}^{n-1}(z;{\bar{t}})}\sum_{j=n-1}^{n}\sum_{\rm part}&g({\bar{t}}^{n-1}_{\scriptscriptstyle\rm III},z)\,\alpha_{n-1}({\bar{t}}^{n-1}_{\scriptscriptstyle\rm III})\,\Omega^{L}({\bar{t}}^{n-1}_{\scriptscriptstyle\rm II},{\bar{t}}^{n-1}_{\scriptscriptstyle\rm III}|{\bar{t}}^{n-2},\varnothing)\times\\ &\quad\times T_{-n,j}(z)\cdot{\mathbb{B}}(\{{\bar{t}}^{s}\}_{0}^{n-2},{\bar{t}}^{n-1}_{\scriptscriptstyle\rm II})\,,\end{split}$ (B.2)
where $|{\bar{t}}^{n-1}_{\scriptscriptstyle\rm I}|=\Theta(-i-n)$ and $|{\bar{t}}^{n-1}_{\scriptscriptstyle\rm III}|=\Theta(j-n)$.
These recurrence relations allows to get an inductive proof of the formula (4.21) similarly to the one presented in appendix A. The only differences for the different steps of this inductive proof are in the different ranges for the indices $\ell$ and $k$ that one has to consider, and for which different identities are needed.
Some of these identities will be different from those used in the proof of the recurrence relations for $\mathfrak{gl}_{n}$ Bethe vectors because of the following mechanism. To consider the inductive step $k+1\to k$ or $\ell\to\ell+1$ in the recurrence relation (4.21) we will apply the zero mode operator ${\sf T}_{k+1,k}$ or ${\sf T}_{\ell+1,\ell}$ to the inductive assumption recurrence relations which correspond to the indices $k+1$ and $\ell$ respectively. Since the action of the zero mode operator ${\sf T}_{k+1,k}$ also parts the set of Bethe parameters ${\bar{t}}^{k}\vdash\{{\bar{t}}^{k}_{\scriptscriptstyle\rm i},{\bar{t}}^{k}_{\scriptscriptstyle\rm ii}\}$, the partition of ${\bar{t}}^{k}$ first by the induction assumption and then by the zero mode action, or vice-versa, may lead to different splittings in the left and right hand sides of the resulting recurrence relation. In the l.h.s. the set ${\bar{t}}^{k}$ first parts into subsets $\{{\bar{t}}^{k}_{\scriptscriptstyle\rm i},{\bar{t}}^{k}_{\scriptscriptstyle\rm ii}\}$ through the action of the zero mode, and then the subset ${\bar{t}}^{k}_{\scriptscriptstyle\rm ii}$ is parted into subsets ${\bar{t}}^{k}_{\scriptscriptstyle\rm ii}\vdash\{{\bar{t}}^{k}_{\scriptscriptstyle\rm I},{\bar{t}}^{k}_{\scriptscriptstyle\rm II}\}$ according to the induction assumption. On the other hand, in the r.h.s. the set ${\bar{t}}^{k}$ first parts into subsets $\{{\bar{t}}^{k}_{\scriptscriptstyle\rm I},{\bar{t}}^{k}_{\scriptscriptstyle\rm II}\}$ through the induction assumption, and then the subset ${\bar{t}}^{k}_{\scriptscriptstyle\rm II}$ is parted into subsets ${\bar{t}}^{k}_{\scriptscriptstyle\rm II}\vdash\{{\bar{t}}^{k}_{\scriptscriptstyle\rm i},{\bar{t}}^{k}_{\scriptscriptstyle\rm ii}\}$ by the action of the zero mode. If the value of the index $j$ is such that according to (4.17) the subset ${\bar{t}}^{k}_{\scriptscriptstyle\rm I}$ is not empty then the resulting equality after action of ${\sf T}_{k+1,k}$ should be symmetrized over the subsets ${\bar{t}}^{k}_{\scriptscriptstyle\rm i}$ and ${\bar{t}}^{k}_{\scriptscriptstyle\rm I}$ both having the cardinality 1. It will make appear the subset ${\bar{t}}^{k}_{\scriptscriptstyle\rm i}\cup{\bar{t}}^{k}_{\scriptscriptstyle\rm I}$, which may have cardinality 2, hence the cardinality 2 subset ${\bar{t}}^{k}_{\scriptscriptstyle\rm I}$ in the sum over partitions in the recurrence relation (4.21). This phenomena does not happen for $\mathfrak{gl}_{n}$ Bethe vectors but is present for the inductive proof of the recurrence relations for the Bethe vectors of other algebra series.
Referring to the calculations presented in appendix A for more details, we describe below the different ranges of the indices $\ell$, $k$ and the corresponding identities which should be used at each of the inductive step. Let us divide the whole domain $-n\leq\ell<k\leq n$ of the values of the indices $\ell$ and $k$ into three subdomains: $1\leq-\ell,k\leq n$, $0\leq\ell<k\leq n$ and $-n\leq\ell<k\leq 0$.
For the first subdomain $1\leq-\ell,k\leq n$, the calculations can be performed as a sequence of the following steps
  • $\mathcal{Z}^{n}_{-n}\cdot{\mathbb{B}}({\bar{t}})\to\mathcal{Z}^{n}_{-n+1}\cdot{\mathbb{B}}({\bar{t}})$ and $\mathcal{Z}^{n}_{-n}\cdot{\mathbb{B}}({\bar{t}})\to\mathcal{Z}^{n-1}_{-n}\cdot{\mathbb{B}}({\bar{t}})$. The recurrence relations for the Bethe vectors ${\mathbb{B}}(\{\bar{w}^{s}\}_{0}^{n-2},\bar{u}^{n-1})=\mathcal{Z}^{n}_{-n+1}\cdot{\mathbb{B}}({\bar{t}})$ and ${\mathbb{B}}(\{\bar{w}^{s}\}_{0}^{n-2},\bar{v}^{n-1})=\mathcal{Z}^{n-1}_{-n}\cdot{\mathbb{B}}({\bar{t}})$ are given by the formulas (B.1) and (B.2) above.
  • $\mathcal{Z}^{n}_{-n+1}\cdot{\mathbb{B}}({\bar{t}})\to\mathcal{Z}^{n}_{\ell}\cdot{\mathbb{B}}({\bar{t}})$, $11\leq-\ell\leq n-$ 1. The recurrence relations (4.21) is proved in these cases by induction over $\ell$ starting from the recurrence relation (B.1). In this case we will not need to perform the symmetrization described above and the only identity which will be necessary to consider these cases is the simple identity (A.6).
  • $\mathcal{Z}^{n}_{\ell}\cdot{\mathbb{B}}({\bar{t}})\to\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})$, $-\ell\leq k\leq n-1$. These cases are proved by induction over $k$ starting from the recurrence relation for the Bethe vector $\mathcal{Z}^{n}_{\ell}\cdot{\mathbb{B}}({\bar{t}})$ which is already proved at the previous step. This proof will require to use the simple identity (A.6) and a more complicated identity
    $\mathop{\rm Sym}_{y_{1},y_{2}}\left(\frac{h(q,y_{1})}{g(y_{1},x)}\ \frac{g(y_{1},y_{2})}{h(y_{2},y_{1})}-\frac{h(y_{1},x)}{g(q,y_{1})}\ \frac{g(y_{2},y_{1})}{h(y_{1},y_{2})}\right)=0,$ (B.3)
    where for any expression $E(y_{1},y_{2})$ we define
    $\mathop{\rm Sym}_{y_{1},y_{2}}E(y_{1},y_{2})=E(y_{1},y_{2})+E(y_{2},y_{1})$
  • $\mathcal{Z}^{n-1}_{-n}\cdot{\mathbb{B}}({\bar{t}})\to\mathcal{Z}^{k}_{-n}\cdot{\mathbb{B}}({\bar{t}})$, $11\leq k\leq n-$ 1. The induction proof of the recurrence relations will require only the simple identity (A.6).
  • $\mathcal{Z}^{k}_{-n}\cdot{\mathbb{B}}({\bar{t}})\to\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})$, $k\leq-\ell\leq n-1$. For these cases the starting point will be the recurrence relation for the Bethe vector $\mathcal{Z}^{k}_{-n}\cdot{\mathbb{B}}({\bar{t}})$ proved at the previous step. We will need for these cases the identities (A.6) and (B.3).
For the second domain $0\leq l<k\leq n$, the most simple way to prove the recurrence relations (4.21) can be depicted as the sequence of the following steps.
  • The first step of the induction $\mathcal{Z}^{n}_{-1}\cdot{\mathbb{B}}({\bar{t}})\to\mathcal{Z}^{n}_{0}\cdot{\mathbb{B}}({\bar{t}})$ is simple. It will not require any complicated rational functions identities, but fixes the function $\psi_{0}(z;{\bar{t}})$ in (4.13).
  • The next step $\mathcal{Z}^{n}_{0}\cdot{\mathbb{B}}({\bar{t}})\to\mathcal{Z}^{n}_{1}\cdot{\mathbb{B}}({\bar{t}})$ is also particular. It is the first step for which partitions with $|\bar{t}^{0}_{\scriptscriptstyle\rm I}|=2$ appear. To get the recurrence relation we use the identities
    $f(t,z)-\frac{f(z,t)}{f(z_{1},t)}=-g(z_{0},t)$ (B.4)
    and
    $\mathop{\rm Sym}_{y_{1},y_{2}}\left(\frac{g(y_{2},z_{0})}{g(x,y_{1})}\Big{(}\mathfrak{f}(y_{1},y_{2})\ f(x,y_{1})\ \frac{f(z,y_{1})}{f(z_{1},y_{1})}-\mathfrak{f}(y_{2},y_{1})\Big{)}\right)=g(z,\bar{y})\,h(x,z)\,,$ (B.5)
    where $\bar{y}=\{y_{1},y_{2}\}$.
  • Then, to prove $\mathcal{Z}^{n}_{\ell}\cdot{\mathbb{B}}({\bar{t}})$, $\ell>1$, starting from $\mathcal{Z}^{n}_{1}\cdot{\mathbb{B}}({\bar{t}})$, we use the identities
    $g(z,\bar{x})\left(h(t,\bar{x})-\frac{f(t,z)}{g(t,\bar{x})}-h(t,z)\right)=g(z,t)$ (B.6)
    and
    $\mathop{\rm Sym}_{y_{1},y_{2}}\left(h(y_{2},z)\left(\frac{h(y_{1},\bar{x})}{g(q,y_{1})}\ \frac{g(y_{2},y_{1})}{h(y_{1},y_{2})}-f(y_{1},z)\frac{h(q,y_{1})}{g(y_{1},\bar{x})}\ \frac{g(y_{1},y_{2})}{h(y_{2},y_{1})}\right)\right)=\frac{g(z,\bar{y})h(q,z)}{g(z,\bar{x})}\,,$ (B.7)
    where $\bar{x}=\{x_{1},x_{2}\}$ and $\bar{y}=\{y_{1},y_{2}\}$.
  • For the steps $\mathcal{Z}^{n}_{\ell}\cdot{\mathbb{B}}({\bar{t}})\to\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})$, $0\leq\ell<k\leq n$, besides the simple identity (A.6), the identity (B.3) should be used.
Finally, for the third domain $-n\leq l<k\leq 0$, the recurrence relations (4.21) is proved in several steps, involving the two particular cases corresponding to $k=0$ and $k=-1$. Among the recurrence relations corresponding to this domain there are the so called shifted recurrence relations, when the sets of Bethe parameters ${\bar{t}}^{s}$ are extended by the shifted parameter $z_{s}=z-c(s-1/2)$.
  • $\mathcal{Z}^{1}_{-n}\cdot{\mathbb{B}}({\bar{t}})\to\mathcal{Z}^{0}_{-n}\cdot{\mathbb{B}}({\bar{t}})$. This step is simple. No complicated identities should be used, but it fixes the function $\phi_{0}(z;{\bar{t}})$ in (4.14).
  • $\mathcal{Z}^{0}_{-n}\cdot{\mathbb{B}}({\bar{t}})\to\mathcal{Z}^{-1}_{-n}\cdot{\mathbb{B}}({\bar{t}})$. Here the proof relies on the identities (B.4) and an identity equivalent to (B.5).
  • For $\mathcal{Z}^{-1}_{-n}\cdot{\mathbb{B}}({\bar{t}})\to\mathcal{Z}^{k}_{-n}\cdot{\mathbb{B}}({\bar{t}})$, $-n<k<-1$, we need identities which appear to be equivalent to (B.6) and (B.7).
  • $\mathcal{Z}^{k}_{-n}\cdot{\mathbb{B}}({\bar{t}})\to\mathcal{Z}^{k}_{\ell}\cdot{\mathbb{B}}({\bar{t}})$, $-n\leq\ell<k\leq 0$. This final step will require identities (B.3).
$\square$